### v1

# CMoE: Fast Carving of Mixture-of-Experts for Efficient LLM Inference

Abstract
Large language models (LLMs) achieve impressive performance by scaling model parameters, but this comes with significant inference overhead. Feed-forward networks (FFNs), which dominate LLM parameters, exhibit high activation sparsity in hidden neurons. To exploit this, researchers have proposed using a mixture-of-experts (MoE) architecture, where only a subset of parameters is activated. However, existing approaches often require extensive training data and resources, limiting their practicality. We propose CMoE (Carved MoE), a novel framework to efficiently carve MoE models from dense models. CMoE achieves remarkable performance through efficient expert grouping and lightweight adaptation. First, neurons are grouped into shared and routed experts based on activation rates. Next, we construct a routing mechanism without training from scratch, incorporating a differentiable routing process and load balancing. Using modest data, CMoE produces a well-designed, usable MoE from a 7B dense model within five minutes. With lightweight fine-tuning, it achieves high-performance recovery in under an hour. We make our code publicly available at https://github.com/JarvisPei/CMoE.
I Introduction
Large language models (LLMs) have demonstrated exceptional proficiency in managing complex tasks and exhibiting emergent capabilities across diverse domains and applications, particularly when scaled to billions of parameters [1, 2, 3, 4]. While LLMs have attained remarkable success, their expanding computational demands and model sizes have intensified challenges related to practical deployment, especially in environments with limited hardware resources or stringent latency requirements. To mitigate these challenges, mixture-of-experts (MoE) architectures [5, 6, 7, 8] have emerged as a promising paradigm. Unlike dense LLMs, where all parameters are activated for every input token, MoE models replace monolithic feed-forward networks (FFNs) with sparsely activated experts: specialized sub-networks that process inputs conditionally via dynamic routing. This design decouples model capacity from computational cost—activating only a subset of experts per token while preserving the model’s expressive power.
Recently, researchers have found that there is high activation sparsity in the hidden neurons of FFNs of dense LLMs, which motivates them to develop sparsity-aware acceleration techniques to reduce computational overhead while maintaining model performance [9, 10]. Building upon this insight, a growing body of research has focused on transforming dense LLMs into MoE architectures through strategic reorganization of FFN parameters but not training MoE from scratch [11, 12, 13]. The prevailing methodology replaces conventional FFN layers with MoE layers, where neurons are partitioned into multiple expert sub-networks while maintaining the original parameter count. During inference, a routing mechanism selectively activates only a subset of experts per input token, thereby achieving dynamic sparsity without compromising model capacity. However, due to the high sparsity target always required in MoE models, these works often need massive computing resources and billions of training data for continual pre-training on the constructed MoE models.
To address the aforementioned limitations, we propose carved MoE, named CMoE, a framework that efficiently carves sparse MoE architectures from dense LLMs through parameter reorganization and training-free structural adaptation. Unlike prior approaches that rebuild MoE models from scratch via resource-intensive pre-training, CMoE strategically “carves” experts from the dense model’s existing feed-forward network (FFN) neurons while preserving their inherent knowledge. This carving process is both computationally efficient—requiring only minutes of lightweight processing—and sophisticated, as it leverages systematic neuron grouping and analytical router construction to retain performance with minimal fine-tuning. The core innovation lies in CMoE’s ability to restructure dense FFNs into MoE layers without re-training the base model. First, we identify neurons that universally encode common knowledge (a.k.a. shared experts) and those that exhibit specialized, input-dependent activation patterns (a.k.a. routed experts). Shared experts are systematically retained by selecting neurons with the highest activation rates, ensuring they capture broadly applicable features. By formulating routed expert grouping as a balanced linear assignment problem, solved via the Jonker-Volgenant algorithm, CMoE clusters neurons into experts while maintaining parameter balance and activation coherence. Second, we derive the routing mechanism directly from the dense model’s activation statistics, bypassing the need for end-to-end router training. This involves constructing a differentiable routing function initialized using representative neurons from each expert cluster, enabling immediate usability while preserving optimization potential.
In summary, the key contributions of this paper are:
  • CMoE: A framework that efficiently carves MoE from dense LLMs by reorganizing FFN neurons into shared/routed experts, eliminating costly pre-training.
  • Shared and routed experts: ensuring parameter balance and activation coherence for CMoE.
  • Training-Free Routing: using representative neurons and lightweight adaptation for router initialization, which enables rapid performance recovery.
Extensive experiments show that, with activation ratio $25\%$, CMoE can maintain a reasonable perplexity even without fine-tuning, and achieve $76.59\%$ accuracies of dense models on some downstream benchmarks with lightweight fine-tuning on 2,048 samples.
II Related Work
In contrast to pretraining MoE models from scratch, recent research has investigated the feasibility of constructing MoE architectures by repurposing existing dense LLMs. Current methodologies for deriving MoE models from dense checkpoints generally follow two paradigms: (1) partitioning parameters of FFNs while preserving the original model’s total parameter count [14, 10], or (2) expanding the model’s overall capacity while retaining activation dimensions comparable to standard dense models [15, 16]. This work prioritizes the former approach. Notably, MoEBERT [14] introduces an importance-driven strategy to transform FFNs into expert modules by strategically redistributing top-scoring neurons across specialized components. Concurrently, MoEfication [10] leverages the discovery of sparse activation patterns in ReLU-based FFNs within T5 architectures, enabling the decomposition of these layers into distinct expert groups governed by a learned routing mechanism. Based on continual training, the LLaMA-2 7B model is modified as a LLaMA-MoE-3.5B MoE model, where the parameters of the original FFNs are partitioned into multiple experts [11]. After training with 200B tokens, the LLaMA-MoE-3.5B model significantly outperforms dense models that contain similar activation parameters. Furthermore, based on a two-stage post-training strategy, an MoE model is constructed from the LLaMA3 8B model, where both attention and MLP are partitioned into MoE blocks [12]].
Extensive experiments have shown the effectiveness of constructing an MoE model from a dense model, and many techniques can be utilized to guarantee performance recovery. However, such performance recovery is extremely resource-consuming, which is unfavorable for efficient deployment in industrial applications. Therefore, more lightweight methods are required, such that performance recovery can be done within hours and even training-free.
Note that model compression such as pruning and quantization is another important technique for efficient LLM inference [17, 18, 19, 20]. Pruning is among the most widely utilized approaches to detect and remove redundant or less significant parameters from models, thereby resulting in a sparser weight matrix and faster inference. ShortGPT [21] has put forward a simple layer-removal approach. This approach is based on block influence, which is determined by the similarity between a layer’s input and output. SliceGPT [22] substitutes each weight matrix with a smaller dense matrix, thereby decreasing the embedding dimension of the network. By taking into account contextual sparsity in real time, Deja Vu [9] and FuseGPT [23] have been proposed to accelerate LLM inference. In contrast to the post-training methods, Learn-To-be-Efficient is designed to train efficiency-aware LLMs so that they learn to activate fewer neurons and achieve a more favorable trade-off between sparsity and performance [24].
III Background
This study primarily focuses on the LLaMA family [2, 25], which uses SwiGLU [26] as the activation function. However, our analysis and findings can be adapted to most of the FFN structures of existing LLMs, including the ReLU-based FFNs [27].
An FFN exists in the tail of each transformer block, which gets the input embedding $\mathbf{x}\in\mathbb{R}^{d}$ and then contributes to the output together with the residual connection, i.e. $\mathbf{x}+F(\mathbf{x})$. Typically, an FFN is a two-layer fully connected network, i.e. the up projection and down projection layer, with an activation function between them. For LLaMA, the SwiGLU composes another gate projection layer. Given the up projection weight $\mathbf{W}_{up}\in\mathbb{R}^{d\times d_{h}}$, the gate projection weight $\mathbf{W}_{gate}\in\mathbb{R}^{d\times d_{h}}$ and the down projection weight $\mathbf{W}_{down}\in\mathbb{R}^{d_{h}\times d}$, the process of an FFN is given by:
$F(\mathbf{x})=\mathbf{h}\mathbf{W}_{down},$ (1)
$\mathbf{h}\!=\!\text{SwiGLU}(\mathbf{x}\mathbf{W}_{up})\!=\!\text{Swish}(\mathbf{x}\mathbf{W}_{gate})\odot(\mathbf{x}\mathbf{W}_{up}),$
where $\text{Swish}(x)=x\cdot\sigma(x)$ is element-wise and $\sigma(\cdot)$ is the sigmoid function.
The basic MoE architecture is composed of a set of $N$ independent FFNs as experts, $\{E_{1},E_{2},...,E_{N}\}$, and a router network $G$ [5]. The output of an MoE-version FFN is then obtained by
$\!\!\!\!F_{MoE}(\mathbf{x})$ $=\!\sum^{N}_{i=1}g_{i}E_{i}(\textbf{x}),$ (2)
$g_{i}$ =$=\!\left\{\begin{array}[]{rl}s_{i},\!&s_{i}\in\text{TopK}(\{s_{i}|1\leq j\leq N\},K),\\ 0,\!&\text{otherwise},\end{array}\right.$
s $=\!\left[s_{1},s_{2},\cdots,s_{N}\right]=G(\textbf{x}),$
where $g_{i}$ is the score for the $i$-th expert, $\textbf{s}\in\mathbb{R}^{N}$ is the token-to-expert affinity, i.e. the output of $G$, and $\text{TopK}(\cdot,K)$ denotes the set comprising $K$ highest scores among the affinity scores calculated for $\mathbf{x}$ on all experts.
IV Methodology
[…]
Figure 1: The overview of our proposed CMoE.
IV-A Shared and Routed Experts Grouping
CMoE starts with grouping neurons in FFN into experts. Shared experts are expected to process common knowledge instead of specialization. Thus, the key idea is to group neurons that are always activated during FFN inference. For routed expert grouping, as described in previous works [28], the key idea is to cluster the neurons that are always activated simultaneously. CMoE carries forward these ideas but comes up with a detailed and analytical perspective. In this section, we construct $N$ experts of size $m=\frac{d_{h}}{N}$ (assuming $d_{h}$ is a multiple of $N$, i.e., $N\mid d_{h}$), including $N_{s}$ shared experts and $N_{r}$ routed experts $(N_{s}+N_{r}=N)$.
We begin with analyzing the isolated contribution of each neuron to the output of FFN. Consider the $i$-th element of $\mathbf{h}\in\mathbb{R}^{d_{h}}$: $h_{i}=\text{Swish}(\mathbf{x}\cdot\mathbf{w}_{gate,i})\cdot(\mathbf{x}\cdot\mathbf{w}_{up,i})$, where $\mathbf{w}_{gate,i}\in\mathbb{R}^{d}$ and $\mathbf{w}_{up,i}\in\mathbb{R}^{d}$ are the $i$-th column of $\mathbf{W}_{gate}$ and $\mathbf{W}_{up}$, respectively. It shows that the value of $h_{i}$ only depends on the input $\mathbf{x}$ and the corresponding $i$-th weights, which implies the independence between neurons. On the other hand, the output of FFN can be written as:
$F(\mathbf{x})=\sum_{i=1}^{d_{h}}h_{i}\mathbf{w}_{down,i},$ (3)
where $\mathbf{w}_{down,i}\in\mathbb{R}^{d}$ is the $i$-th row of $\mathbf{W}_{down}$. When we revisit the FFN process, we can regard each $h_{i}$ as the score to be multiplied to the split vector $\mathbf{w}_{down,i}$, whose product contributes to part of the output $F(\mathbf{x})$. As a finding of some structured pruning research [29, 30], the norm of $F(\mathbf{x})$ is always small due to the residual connection. Such phenomenon implies the high sparsity existed in FFN, and by (3) we relate it to $h_{i}$, whose value decides how large an isolated neuron contributes to the output. Therefore we make a hypothesis:
$\arg\min_{i}|h_{i}\mathbf{w}_{down,i}|\approx\arg\min_{i}|h_{i}|,$ (4)
which is reasonable since when $h_{i}$ is extremely small, the product $h_{i}\mathbf{w}_{down,i}$ will also vanish. It is expected that the hidden state $\mathbf{h}$ should be highly sparse, which means $|h_{i}|$ is often extremely small. To verify it, we hack into the FFN and draw the distribution of $\mathbf{h}$.
Figure 2: The histogram of FFN hidden state $\mathbf{h}$ for the $3$-th block and the $1,000$-th token.
As demonstrated in fig. 2, the distribution is sharply peaked at 00 and constrained within a small range, indicating that most $h_{i}$ are concentrated near zero and confirming the sparsity. And it exhibits symmetry, suggesting a near-normal distribution centered around the mean. Based on what we discuss and observe above, the hidden state values work well as the basis for judgment of neuron activation, because of its high differentiation and independence across different neurons. Therefore, we propose a new metric, called absolute TopK (ATopK), to determine the activation status of a neuron with index $i$:
$a_{i}=\left\{\begin{array}[]{rl}1,&\lvert h_{i}\rvert\in\text{TopK}(\{\lvert h_{j}\rvert\ |1\leq j\leq d_{h}\},K_{a}),\\ 0,&\text{otherwise},\end{array}\right.$ (5)
where we choose neurons with $K_{a}$ highest absolute value among the hidden state values, and assign their labels in the activation marker $\mathbf{a}=\left[a_{1},a_{2},\cdots,a_{d_{h}}\right]$ with $1$.
To further evaluate the activation information, we make samples in the training set to record their activation markers. Given a batched input tensor $\mathbf{X}\in\mathbb{R}^{b\times s\times d}$, with $b$ the batch size and $s$ the sequence length, we obtain the batched hidden state $\mathbf{H}\in\mathbb{R}^{b\times s\times d_{h}}$ as follows:
$\mathbf{H}=\text{Swish}(\mathbf{X}\mathbf{W}_{gate})\odot(\mathbf{X}\mathbf{W}_{up}).$ (6)
Note that in practical implementation, we normalize $\mathbf{X}$, $\mathbf{W}_{gate}$ and $\mathbf{W}_{up}$ before the calculation to eliminate the influence of their magnitudes on the output. We then calculate the activation markers for the hidden state of all the tokens, which is reshaped as an activation feature matrix $\mathbf{A}\in\mathbb{R}^{q\times d_{h}},q=b\cdot s$. Denote the $i$-th column of $\mathbf{A}$ as the feature vector $\mathbf{c}_{i}\in\mathbb{R}^{q}$, i.e., $\mathbf{A}=[\mathbf{c}_{1}\;\mathbf{c}_{2}\;\cdots\;\mathbf{c}_{d_{h}}]$ which represents the $i$-th neuron’s activation status on the sampled $q$ tokens. By calculating the expected value of each $\mathbf{c}_{i}$, we can obtain the vector of activated rates $\bm{\mu}$ as:
$\bm{\mu}=[\mu_{1},\mu_{2},\cdots,\mu_{d_{h}}],\mbox{ with }\mu_{i}=\frac{1}{q}\sum_{j=1}^{q}c_{j,i},$ (7)
where $c_{j,i}$ indicates the $j$-th element of $\mathbf{c}_{i}$. We draw the histogram of $\bm{\mu}$ as in Fig. 3.
Figure 3: The histogram of activation rates $\bm{\mu}$ for the $3$-th block with $K_{a}=1,000$.
The histogram reveals a highly skewed distribution of activation rates, where the majority of neurons exhibit low activation rates (below 0.1), with a sharp peak near 0.07. However, the distribution also features a long tail, indicating the presence of a subset of neurons with significantly higher activation rates extending up to 1. These high-activation neurons are likely active across a wide range of input tokens, making them suitable for processing common knowledge rather than task-specific specialization. Therefore, we construct shared experts by grouping these high-activation neurons. Given the total number of shared experts as $N_{s}$ and the expert size $m$, we get the selection indices set $S_{N_{s}}$ by selecting $N_{s}\cdot m$ neurons with highest activation rates:
$S_{N_{s}}=\{i\ :\ \mu_{i}\in\text{TopK}(\{\mu_{j}\mid 1\leq j\leq d_{h}\},N_{s}\cdot m)\}.$ (8)
Since the independency across neurons, as we discussed, we build the experts using the same architecture as the original FFN and assign the expert parameters to the neurons of selected indices. Since all shared experts are always activated, we just construct them together for convenience:
$\mathbf{W}_{up}^{s}$ $=\mathbf{W}_{up}[:,S_{N_{s}}],$ (9)
$\mathbf{W}_{gate}^{s}$ $=\mathbf{W}_{gate}[:,S_{N_{s}}],$
$\mathbf{W}_{down}^{s}$ $=\mathbf{W}_{down}[S_{N_{s}},:],$
where $\mathbf{W}_{up}^{s}\in\mathbb{R}^{d\times N_{s}\cdot m}$, $\mathbf{W}_{gate}^{s}\in\mathbb{R}^{d\times N_{s}\cdot m}$ and $\mathbf{W}_{down}^{s}\in\mathbb{R}^{N_{s}\cdot m\times d}$ are the weights of linear layers of the shared experts $E^{s}$. Given the input embedding $\mathbf{x}$, the output of $E^{s}$ is obtained via
$E^{s}(\mathbf{x})\!=\!\mathbf{h}^{s}\mathbf{W}_{down}^{s},\mbox{with }\mathbf{h}^{s}\!=\!\text{Swish}(\mathbf{x}\mathbf{W}_{gate}^{s})\!\odot\!(\mathbf{x}\mathbf{W}_{up}^{s}).$
The majority of low activation rates also encourage us to construct routed experts, which are not always activated but are specialized for tokens encountered. To construct $N_{r}$ routed experts, we develop a customized balanced K-means clustering algorithm. We first identify $N_{r}$ centroids as the neurons (excluding those assigned to shared experts) with the highest activation rates, and group their feature vectors into a centroid set $C=\{\mathbf{c}_{i}:\mu_{i}\in\text{TopK}(\{\mu_{j}\mid 1\leq j\leq d_{h},j\notin S_{Ns}\},N_{r})\}=\{\hat{\mathbf{c}}_{1},\dots,\hat{\mathbf{c}}_{N_{r}}\}$, where, for convenience, we re-label the centroids with $\hat{\mathbf{c}}$.
By selecting these centroids, we ensure that the clusters are initialized around meaningful and prominent activation patterns, providing a strong starting point for clustering. Using the centroids selected, we will construct clusters with remaining neurons. We pay attention to their correlation to the centroids, i.e. we study the distances of their feature vectors to the centroids’. The feature vector $\mathbf{c}$ individually contains the activation status of a specific neuron, and we expect that neurons with similar functionalities will have similar patterns on the activation status, i.e. they are often activated and deactivated together during inference. Therefore, we compute the $L_{2}$ distance of each feature vector $\mathbf{c}$ to each centroid $\hat{\mathbf{c}}$. Define a distance matrix $\mathbf{D}\in\mathbb{R}^{N_{r}\cdot m\times N_{r}}$, whose element $d_{i,j}$ is the $L_{2}$ distance between the $i$-th feature vector $\mathbf{c}_{i}$ and the $j$-th centroid $\hat{\mathbf{c}}_{j}$:
$d_{i,j}=\|\mathbf{c}_{i}-\hat{\mathbf{c}}_{j}\|_{2}=\sqrt{\sum_{k=1}^{q}(c_{k,i}-\hat{c}_{k,j})^{2}},$ (10)
where, for convenience, we also include the original feature vectors of centroids inside the set to be grouped.
To group the neurons into routed experts, with the centroids designed above at $t=0$, the constrained balanced $K$-means clustering algorithm works as follows [31]. Given $N_{r}\cdot m$ vectors $\mathbf{c}$, the cluster membership value $m$ and cluster centroids $\hat{\mathbf{c}}_{1}^{t},\hat{\mathbf{c}}_{2}^{t},\dots,\hat{\mathbf{c}}_{N_{r}}^{t}$ at iteration $t$, compute $\hat{\mathbf{c}}_{1}^{t+1},\hat{\mathbf{c}}_{2}^{t+1},\dots,\hat{\mathbf{c}}_{N_{r}}^{t+1}$ at iteration $t+1$ with the following $2$ steps:
Cluster Assignment Let $T_{i,p}^{t}$ be a solution to the following linear program with $\hat{\mathbf{c}}_{p}^{t}$ fixed:
$\min_{T}$ $\sum_{i=1}^{N_{r}\cdot m}\sum_{p=1}^{N_{r}}T_{i,p}\cdot d_{i,p}$ (11)
s.t. $\sum_{i=1}^{N_{r}\cdot m}T_{i,p}=m,p=1,\dots N_{r}$
$\sum_{p=1}^{N_{r}}T_{i,p},i=1,\dots,N_{r}\cdot m$
$T_{i,p}\geq 0,i=1,\dots,N_{r}\cdot m,p=1,\dots N_{r}.$
Cluster Update Update $\hat{\mathbf{c}}_{p}^{t+1}$ as follows:
$\hat{\mathbf{c}}_{p}^{t+1}=\left\{\begin{array}[]{ll}\frac{\sum_{i=1}^{N_{r}\cdot m}T_{i,p}^{t}\cdot\mathbf{c}_{i}}{\sum_{i=1}^{N_{r}\cdot m}T_{i,p}^{t}},&\text{if}\sum_{i=1}^{N_{r}\cdot m}T_{i,p}^{t}>0,\\ \hat{\mathbf{c}}_{p}^{t},&\text{otherwise}.\end{array}\right.$ (12)
The steps are terminated until $\hat{\mathbf{c}}_{p}^{t+1}=\hat{\mathbf{c}}_{p}^{t}$, $\forall p=1,\dots,N_{r}$.
We include the distance value $d_{i,j}$ into the cost function so that the balanced $K$-means clustering makes the intra-expert distances low and inter-expert distances high. However, the cluster assignment we defined above is an unbalanced assignment problem ($N_{r}\cdot m>m$) and cannot directly be solved by existing algorithms for balanced assignment. Therefore, we reduce it to a balanced assignment by extending the distance matrix $\mathbf{D}$ as follows:
$\mathbf{D}^{ext}$ $=\big{[}\underbrace{\mathbf{d}_{1},\dots,\mathbf{d}_{1}}_{m\text{times}},\underbrace{\mathbf{d}_{2},\dots,\mathbf{d}_{2}}_{m\text{ times}},\dots,\underbrace{\mathbf{d}_{N_{r}},\dots,\mathbf{d}_{N_{r}}}_{m\text{ times}}\big{]},$
where we repeat every column $\mathbf{d}$ of $\mathbf{D}$ $m$-times to obtain the extended matrix $\mathbf{D}^{ext}\in\mathbb{R}^{N_{r}\cdot m\times N_{r}\cdot m}$. Then the reduced balanced assignment problem is formulated as:
Balanced Assignment Let $T_{i,p^{\prime}}^{{}^{\prime},t}$ be a solution to the following linear program with $\hat{\mathbf{c}}_{p}^{t}$ fixed:
$\min_{T^{\prime}}$ $\sum_{i=1}^{N_{r}\cdot m}\sum_{p^{\prime}=1}^{N_{r}\cdot m}T^{\prime}_{i,p^{\prime}}\cdot d_{i,p^{\prime}}^{ext}$ (13)
s.t. $\sum_{i=1}^{N_{r}\cdot m}T^{\prime}_{i,p^{\prime}}=1,p^{\prime}=1,\dots N_{r}\cdot m$
$\sum_{p^{\prime}=1}^{N_{r}\cdot m}T^{\prime}_{i,p^{\prime}}=1,i=1,\dots,N_{r}\cdot m$
$T^{\prime}_{i,p^{\prime}}\geq 0,i=1,\dots,N_{r}\cdot m,p^{\prime}=1,\dots N_{r}\cdot m,$
where $d_{i,p^{\prime}}^{ext}=d_{i,p}$ if $p\cdot m<p^{\prime}\leq(p+1)\cdot m$.
And the balanced cluster $\hat{\mathbf{c}}_{p}^{t+1}$ can be updated as follows:
$\hat{\mathbf{c}}_{p}^{t+1}=\scriptstyle\left\{\begin{array}[]{ll}\frac{\sum\limits_{k=k_{0}}^{mp}\sum\limits_{i=1}^{N_{r}\cdot m}T_{i,k}^{{}^{\prime},t}\cdot\mathbf{c}_{i}}{\sum\limits_{k=k_{0}}^{mp}\sum\limits_{i=1}^{N_{r}\cdot m}T_{i,k}^{{}^{\prime},t}},&\text{if}\scriptstyle\sum\limits_{k=k_{0}}^{mp}\sum\limits_{i=1}^{N_{r}\cdot m}T_{i,k}^{{}^{\prime},t}>0,\\ \hat{\mathbf{c}}_{p}^{t},&\text{otherwise},\end{array}\right.$ (14)
where $k_{0}=m(p-1)+1$. Drawing on the Jonker-Volgenant algorithm [32], this problem can be addressed as a reduced assignment problem in each step of the $K$-means algorithm, with a complexity of $O(n^{3})$.
The final solution $T^{\prime}$ gives us the optimized strategy to group the routed experts. We get the selection indices set $S_{N_{r},p}$, $p=1,\dots,N_{r}$, for each routed expert $E^{r}_{p}$:
$S_{N_{r},p}=\{i\ :\exists\ T^{\prime}_{i,k}=1,\ \text{for}\ \forall k\in\{m(p-1)+1,\dots,m\}\}.$
We then build the weights $\mathbf{W}_{up}^{r,p}\in\mathbb{R}^{d\times m}$, $\mathbf{W}_{gate}^{r,p}\in\mathbb{R}^{d\times m}$ and $\mathbf{W}_{down}^{r,p}\in\mathbb{R}^{m\times d}$ of each routed expert $E_{p}^{r}$ as in Equation 9. And the output of $E_{p}^{r}$ is given by:
$E^{r}_{p}(\mathbf{x})\!=\!\mathbf{h}^{r}_{p}\mathbf{W}_{down}^{r,p},\mbox{ with }\mathbf{h}^{r}_{p}\!=\!\text{Swish}(\mathbf{x}\mathbf{W}_{gate}^{r,p})\odot(\mathbf{x}\mathbf{W}_{up}^{r,p}).$
Suppose the MoE activates $N_{k}$ experts out of the $N_{r}$ routed experts. We expect the router to choose the routed experts with top $N_{k}$ scores. Therefore, we modify the MoE-version FFN as in (IV-A),
$F_{MoE}(\mathbf{x})$ $=E^{s}(\mathbf{x})+\sum^{N_{r}}_{i=1}g_{i}E_{i}^{r}(\textbf{x}),$
$g_{i}$ =$=\left\{\begin{array}[]{rl}1,&s_{i}\in\text{TopK}(\{s_{i}|1\leq j\leq N_{r}\},N_{k}),\\ 0,&\text{otherwise},\end{array}\right.$ (17)
s $=\left[s_{1},s_{2},\cdots,s_{N_{r}}\right]=G(\textbf{x}),$ (18)
where we make the expert score $g_{i}\in\{0,1\}$ to enable an un-scaled version of the expert output to avoid biases.
IV-B Training-free Router Construction
We now present the training-free router network $G$ for CMoE. Unlike previous works, which either built the router from scratch or intuitively used hidden features as initialization, we formulate the router construction as a minimization problem and then develop an algorithm with analysis to construct the router by approximating the optimal solution.
Given the same input embedding $\mathbf{x}$, the output of original output of the dense FFN, i.e. $F(\mathbf{x})$ in (1), is equivalent to the sum of the output of all the experts in $F_{MoE}$:
$F(\mathbf{x})$ $=E^{s}(\mathbf{x})+\sum^{N_{r}}_{i=1}E_{i}^{r}(\textbf{x}).$ (19)
The only difference between (19) and $F_{MoE}(\mathbf{x})$ in (IV-A) is the expert score $\mathbf{g}$, which is obtained from the TopK selection of the output of $G$. Therefore, to preserve important knowledge captured by the original dense FNN, $G$ can be constructed to enforce $F_{MoE}(\mathbf{x})$ to be close to $F(\mathbf{x})$ by solving the minimization problem (IV-B),
$\arg\min_{G}|F_{MoE}(x;G)-F(x)|$
=\!= $\!\arg\min_{G}|\sum^{N_{r}}_{i=1}(g_{i}-1)E_{i}^{r}(\textbf{x})|=\!\!\arg\min_{G}|\sum^{i\in S_{de}}E_{i}^{r}(\mathbf{x})|,$
where $S_{de}\!=\!\{i:s_{i}\!\notin\!\text{TopK}(\{s_{i}|1\!\leq\!j\!\leq\!N_{r}\},N_{k})\}$ and $|S_{de}|\!=\!N_{r}\!-\!N_{k}$, and the problem becomes constructing the $G$ to minimize the absolute sum of the output of deactivated routed experts. Note that we have made a hypothesis in (4) that the output/sparsity of $F(x)$ is highly related to the norm/sparsity of $\mathbf{h}$, which is the same for the expert outputs. Based on (3) and (4), we reformulate (IV-B) as in (20):
$\arg\min_{G}|\sum^{i\in S_{de}}E_{i}^{r}(\mathbf{x})|$ (20)
[…]
=\!= $\!\arg\min_{G}\mathbb{E}_{\mathbf{h}}\left[\|\mathbf{h}^{r}_{i}\|_{1}\mid i\in S_{de}\right].$
The problem becomes constructing $G$ that can minimize the expected hidden states $\mathbf{h}$ of deactivated routed experts. Note that $G$ controls the de-/activation of routed experts by outputting the token-to-expert affinity $\mathbf{s}=[s_{1},\dots,s_{N_{r}}]$. Therefore, a solution for the above problem is to construct the $G$ such that matching the sorting indices of the set $\{s_{1},\dots,s_{N_{r}}\}$ and the set $\{\bar{\mathbf{h}}^{r}_{1},\dots,\bar{\mathbf{h}}^{r}_{N_{r}}\}$ ($\bar{\mathbf{h}}^{r}_{i}=\mathbb{E}_{\mathbf{h}}\left[||\mathbf{h}^{r}_{i}||_{1}\right]$), i.e. $\exists\ \text{permutation}\ \sigma$ such that:
$s_{\sigma(1)}\leq s_{\sigma(2)}\leq\dots\leq s_{\sigma(N_{r})}\mbox{ and}$
$\bar{\mathbf{h}}^{r}_{\sigma(1)}\leq\bar{\mathbf{h}}^{r}_{\sigma(2)}\leq\dots\leq\bar{\mathbf{h}}^{r}_{\sigma(N_{r})},$ (21)
by which we can verify that the minimum of (20) is:
$\min_{G}$ $\mathbb{E}_{\mathbf{h}}\left[\|\mathbf{h}^{r}_{i}\|_{1}\mid i\in S_{de}\right]=\frac{1}{N_{r}-N_{k}}\!\!\sum_{i=1}^{N_{r}-N_{k}}\bar{\mathbf{h}}^{r}_{\sigma(i)},$ (22)
which is obtained by setting $G(\mathbf{x})$ as:
$S_{de}$ =$=\{\sigma(1),\dots,\sigma(N_{r}-N_{k})\}$
=$=\{i:s_{i}\notin\text{TopK}(\{s_{i}|1\leq j\leq N_{r}\},N_{k})\}.$ (23)
Finally, we pay attention to the hidden state $\mathbf{h}^{r}$ of any routed expert, which is the output by the neurons that we grouped in Section IV-A, where the cluster is centered to the centroid $\hat{\mathbf{c}}$. We denote the neuron in each cluster that has the lowest $L_{2}$ distance to the centroid as the representative neuron:
$R_{j}=i,\mbox{ if }\|\mathbf{c}_{i}-\hat{\mathbf{c}}_{j}\|_{2}\leq\|\mathbf{c}_{k}-\hat{\mathbf{c}}_{j}\|_{2},\forall k\in S_{N_{r},j}.$ (24)
Therefore, when we regard the hidden state value $h^{r}_{R_{j}}$ as feature of the representative neuron and assume that $h^{r}_{R_{j}}\approx\bar{\mathbf{h}}^{r}_{j}$, where $\bar{\mathbf{h}}^{r}_{j}$ refers to the expected hidden state value, we can construct the router by grouping the representative neurons of all the routed experts:
$G(\mathbf{x})$ $=\text{Swish}(\mathbf{x}\mathbf{W}_{gate}^{R})\odot(\mathbf{x}\mathbf{W}_{up}^{R}),$ (25)
where $\mathbf{W}_{gate}^{R}=\mathbf{W}_{gate}[:,S_{R}]$, $\mathbf{W}_{up}^{R}=\mathbf{W}_{up}[:,S_{R}]$, and $S_{R}=\{R_{1},\dots,R_{N_{r}}\}$. This leads to
$G(\mathbf{x})$ $=\left[s_{1},s_{2},\cdots,s_{N_{r}}\right]=\left[h^{r}_{R_{1}},h^{r}_{R_{2}},\cdots,h^{r}_{R_{N_{r}}}\right]$
$\approx\left[\bar{\mathbf{h}}^{r}_{1},\mathbf{h}^{r}_{2},\cdots,\mathbf{h}^{r}_{N_{r}}\right],$ (26)
which is hence an approximate solution for the original problem introduced in (IV-B).
IV-C Differentiable Routing and Load-balancing
Though we have constructed a well-designed router in Section IV-B, it is not differentiable since each expert score $g_{i}$ is a constant, as shown in (IV-A), hindering further alignment and performance recovery. Therefore, we introduce a learnable parameter $\mathbf{u}$ when computing the expert scores as follows,
$g_{i}\!$ =$\!=\!\left\{\begin{array}[]{ll}\!\!1+s^{\prime}_{i}\cdot u_{i},&s^{\prime}_{i}\in\text{TopK}(\{s^{\prime}_{i}|1\leq j\leq N_{r}\},N_{k}),\\ \!\!0,&\text{otherwise},\end{array}\right.$ (28)
$\mathbf{s}^{\prime}$ $=\text{Softmax}(\mathbf{s}),\ \textbf{u}=[u_{1},u_{2},\dots,u_{N_{r}}]$ (29)
where the scale $\mathbf{u}$ is initialized as zero to avoid perturbation.
For MoE models, load-balancing is crucial to guarantee computational efficiency, especially expert parallelism in LLMs serving. As in DeepSeek-V3 [4], we use the auxiliary-loss-free load balancing by introducing a bias term $\mathbf{b}$ before the TopK selection:
$g_{i}\!$ =$\!=\!\left\{\begin{array}[]{ll}\!\!1+s^{\prime}_{i}\cdot u_{i},&s^{\prime}_{i}+b_{i}\in\text{TopK}(\{s^{\prime}_{i}|1\leq j\leq N_{r}\},N_{k}),\\ \!\!0,&\text{otherwise},\end{array}\right.$ (32)
$\mathbf{b}$ $=[b_{1},b_{2},\dots,b_{N_{r}}].$ (33)
Here, $\mathbf{b}$ is initialized as zero and updated based on the expert load status during each step of training, as in DeepSeek-V3. The hyper-parameter update speed $\gamma$ is used to update the bias term $\mathbf{b}$ at the end of each step, i.e. decreasing/increasing the bias term by $\gamma$ if the corresponding expert is overloaded/underloaded.
V Experiments
TABLE I: Comparison of perplexity results. Randomly select samples from WikiText-2 training dataset for calibration and training.
Method Type LLaMA-2-7B LLaMA-3-8B Training-free Fine-tuning Training-free Fine-tuning WikiText-2 C4 WikiText-2 C4 WikiText-2 C4 WikiText-2 C4 Dense - 5.27 7.27 - - 6.14 9.44 - - LLaMA-MoE A2E8 nan nan 468.00 2660.68 nan nan 988.20 7521.83 LLaMA-MoE A4E16 nan nan 540.62 2690.68 nan nan 1094.24 7758.91 CMoE S1A1E8 60.86 135.61 12.76 32.12 162.74 324.71 21.16 64.03 CMoE S1A3E16 89.19 180.65 13.84 33.56 262.85 465.09 22.97 72.34 CMoE S2A2E16 62.30 136.12 12.73 32.37 143.38 284.19 21.01 65.57
CMoE is implemented based on Hugging Face Transformers [33] together with Pytorch [34]. The experiments are conducted on one NVIDIA H800 PCIe 80GB graphics card with CUDA Driver 12.6. We randomly select samples from WikiText-2 training dataset [35] as calibration and fine-tuning data. We use only 8 examples with 2,048 sequence length as calibration for calculating the hidden state $\mathbf{H}$ in (6). We set $K_{a}=10$ for the activation status record, which sounds counter-intuitive but works best in practice. For lightweight fine-tuning, we run $1$ epoch using the Adam optimizer [36] with $\beta_{1}$ = 0.9 and $\beta_{2}$ = 0.95. We also employ LoRA [37] with a rank of 8 and $lora\_alpha=32$. We fine-tune all the models including the baselines with $2,048$ samples. We set different initial learning rates for the score scale $\mathbf{u}$ and other parameters, i.e. $0.001$ and $5.95e^{-5}$, respectively. We set the bias update speed $\gamma$ to $0.001$.
V-A Main Results
We compare CMoE with the up-to-date baseline LLaMA-MoE [11], in which the neurons are randomly split, and continual pre-training is carried out with additional router networks. The experiments cover both training-free and lightweight fine-tuning versions. We design these experiments to demonstrate the remarkable post-training performance of CMoE. we demonstrate the results on two different datasets, i.e. WikiText-2 [35] and C4 [38]. The baseline models are chosen as LLaMA-2-7B and LLaMa-3-8B.
Language Modeling Performance. We evaluate the perplexity of the constructed Mixture-of-Experts (MoE) models in Table I. We use abbreviations to denote the composition of experts. For instance, ‘S2A2E16’ represents 2 shared experts, 2 activated routed experts, and a total of 16 experts and the total activation ratio is $(2+2)/16=25\%$. Our findings indicate that in the training-free scenario, the ppl of our baseline LLaMA-MoE on both LLaMA-2-7B and LLaMA-3-8B are NaN across different datasets. In contrast, the perplexity of the proposed CMoE can be effectively controlled. Moreover, after fine-tuning, we observe that the perplexity can be reduced to as low as 12.73 when the activation ratio is 25%, which corresponds to a sparsity of 75%. Additionally, the CMoE model outperforms all state-of-the-art (SOTA) structured pruning methods [9, 22, 23].
Downstream Tasks Performance. We also evaluate the performance of the constructed MoE models on various downstream tasks. We evaluate on the following benchmarks: 32-shot BoolQ [39], 0-shot PIQA [40], 0-shot SciQ [41], 5-shot Winogrande [42], 25-shot ARC-Challenge [43], and 10-shot HellaSwag [44]. The results are presented in table II, where we denote fine-tuning or not with ‘FT’. In all activation ratio configurations, we find that the proposed CMoE outperforms the LLaMA-MoE in accuracy across a diverse set of downstream tasks, on both training-free and fine-tuning scenarios.
To illustrate the performance of the models, we select the SciQ dataset for zero-shot testing and the BoolQ dataset for few-shot testing as representative examples. When the activation ratio is set at a fixed value of 25% (corresponding to a sparsity rate of 75%), after fine-tuning, the LLaMA-MoE model can only achieve an accuracy equivalent to 21.2% of that of the relevant dense model on the SciQ dataset and 45.31% on the BoolQ dataset. Given that the perplexity (ppl) of the LLaMA-MoE model is Not a Number (NaN) in the training-free scenario, we do not report its accuracy in this case. In contrast, even without fine-tuning (i.e., in the training-free scenario), the proposed CMoE model demonstrates superior performance, attaining an accuracy equivalent to 56.3% of that of the dense model on the SciQ dataset and 54% on the BoolQ dataset. After fine-tuning, the performance of the CMoE model is further enhanced, with the accuracy reaching 76.59% on the SciQ dataset and 74.3% on the BoolQ dataset (relative to dense model).
V-B Ablation Studies
We conduct ablation studies with LLaMA-2-7B and data randomly selected from WikiText-2 training datasets.
Figure 4: Trade-off between Model Performance and Construction Time with Increasing Training Data.
TABLE II: Comparison of downstream tasks. Randomly select samples from WikiText-2 training dataset as calibration and training data.
Model Method Type FT BoolQ(32) SciQ PIQA WinoGrande(5) ARC-C(25) HellaSwag(10) LLaMA-2-7B Dense - - 82.04 90.80 78.78 73.95 53.15 78.55 LLama-MoE A2E8 37.83 20.00 49.73 50.12 25.79 26.18 LLama-MoE A4E16 37.82 20.60 49.56 49.40 25.98 27.21 CMoE S1A1E8 46.09 65.30 52.77 48.70 23.80 30.12 CMoE S1A3E16 50.79 62.00 52.29 50.12 24.23 28.76 CMoE S2A2E16 53.85 66.50 53.26 49.33 23.72 29.85 CMoE S1A1E8 55.04 77.50 57.12 54.06 27.56 38.79 CMoE S1A3E16 53.66 75.30 56.37 53.98 27.82 38.89 CMoE S2A2E16 57.19 77.30 56.86 52.57 26.45 38.90 LLaMA-3-8B Dense - - 83.48 94.2 80.79 77.50 59.72 82.16 LLama-MoE A2E8 37.82 20.30 49.02 50.74 25.68 25.76 LLama-MoE A4E16 37.83 20.50 51.52 49.96 25.27 26.00 CMoE S1A1E8 47.43 47.70 52.56 50.67 23.72 28.08 CMoE S1A3E16 41.31 46.60 51.63 48.77 24.15 27.41 CMoE S2A2E16 45.08 53.10 50.87 53.19 24.23 28.44 CMoE S1A1E8 56.88 72.00 57.34 51.93 25.77 36.68 CMoE S1A3E16 60.55 71.40 57.07 53.51 26.54 36.28 CMoE S2A2E16 62.14 72.20 59.08 51.14 27.73 36.68
Figure 5: Ablation studies:(a) Impact of shared expert ratio on model performance; (b) Activation Rate vs. Model Performance; (c) Effect of Load Balancing.
Impact of Training Data Size. To quantify the relationship between training data volume, model performance, and computational efficiency, we systematically vary the number of training samples from 0 (initial state) to 4,096 while measuring perplexity (PPL) and construction time. Fig. 4 illustrates the results on the setting of “S2A2E16”, revealing critical trends in performance-cost trade-offs.
Increasing training data from 0 to 4,096 samples reduces perplexity by 80.4% (62.30 $\to$ 12.21) but non-linearly increases construction time from 298s to 4,502s. While performance improves sharply initially (64 samples cut PPL by 63%), gains diminish beyond 1,024 samples (13.47 PPL), with marginal improvements (12.21 PPL at 4,096 samples) requiring disproportionately higher runtime (+133% from 2,048 to 4,096 samples). The results demonstrate that CMoE only needs a modest of data to achieve low perplexity, while further performance increase is hard and possibly demands more diverse data.
Shared Experts Ratio. We carry out an extensive analysis of the influence of the shared expert ratio on perplexity. In our experimental setup, we have a total of 32 experts, among which 8 are activated. We systematically adjust the proportion of shared experts within these 8 activated experts. As depicted in Table Fig. 5, the results clearly demonstrate a distinct trend: as the ratio of shared experts rises from 0.125 to 0.875, the perplexity continuously decreases. There is a remarkable improvement in the perplexity value, dropping from 14.48 to 11.93. This finding indicates that shared experts generally contribute positively to the performance of the model. However, it is also evident that the marginal returns tend to diminish when the ratio of shared experts reaches a relatively high level.
Activation Rate. We conduct an external analysis to investigate how the total activation rate (comprising both shared and activated routed experts) impacts model performance across different domains, in comparison with dense baselines. We perform evaluations on the CMoE model featuring a total of 16 experts. Specifically, we vary the ratio of shared experts and activated routed experts at a fixed proportion of 1:1 on the WikiText-2 and C4 datasets. As illustrated in Table Fig. 5, a monotonic decrease in PPL can be found as the activation rate increases, gradually approaching the performance level of the dense model. For both datasets, an activation rate of 75% enables the model to achieve nearly comparable performance to the dense model. On the WikiText-2 dataset, the PPL values are 5.79 for the model under consideration and 5.27 for the dense model; on the C4 dataset, the corresponding values are 11.19 and 7.27. These results demonstrate that CMoE architectures can attain a performance quality comparable to that of dense models, even when characterized by relatively high sparsity.
Load Balancing. CMoE can inherently achieve effective load-balancing among routed experts, with the exception of some special blocks. For example, the last block of LLaMA-2-7B exhibits extremely unbalanced expert counts. However, this issue can be effectively resolved by the load-balancing mechanism we introduced in Section IV-C. As presented in Fig. 5, prior to the implementation of the load-balancing strategy, substantial disparities in the workloads assigned to different experts are clearly observable. Specifically, Expert 3 and Expert 8 are burdened with disproportionately high loads, handling 11,163 and 11,225 instances respectively. In stark contrast, other experts process a significantly lower number of instances. For example, Expert 12 and Expert 13 handle only 100 and 102 instances respectively. Upon the application of our proposed load - balancing mechanism, a more equitable and uniform distribution of the computational workload across all experts is successfully achieved. After the adjustment, the number of instances processed by each expert falls within the range of 1,443 to 3,584.
VI Conclusion
We present CMoE, a framework that efficiently carves sparse Mixture-of-Experts (MoE) architectures from dense LLMs through parameter reorganization and lightweight adaptation. By leveraging activation sparsity patterns in FFN layers, CMoE groups neurons into shared and routed experts via a novel balanced linear assignment formulation. The router is analytically initialized from activation statistics and refined via differentiable scaling and load balancing. Experiments show that CMoE can construct effective Mixture-of-Experts (MoE) models, achieving comparable perplexity to dense models and outperforming baseline models. For instance, on the SciQ dataset, CMoE can reach 56.3% even without fine-tuning and it can further improve the accuracy to 76.59% after fine-tuning, a huge outperforming compared with 21.2% (the accuracy of the baseline LLaMA-MoE). Extensive experiments have demonstrated that CMoE offers a practical approach for deploying large language models (LLMs) in resource-constrained environments.

### v2

@@
-CMoE: Fast Carving of Mixture-of-Experts for Efficient LLM Inference
+CMoE: Converting Mixture-of-Experts from Dense to Accelerate LLM Inference
 Abstract
-Large language models (LLMs) achieve impressive performance by scaling model parameters, but this comes with significant inference overhead.
-Feed-forward networks (FFNs), which dominate LLM parameters, exhibit high activation sparsity in hidden neurons.
-To exploit this, researchers have proposed using a mixture-of-experts (MoE) architecture, where only a subset of parameters is activated.
-However, existing approaches often require extensive training data and resources, limiting their practicality.
-We propose CMoE (Carved MoE), a novel framework to efficiently carve MoE models from dense models.
-CMoE achieves remarkable performance through efficient expert grouping and lightweight adaptation.
-First, neurons are grouped into shared and routed experts based on activation rates.
-Next, we construct a routing mechanism without training from scratch, incorporating a differentiable routing process and load balancing.
-Using modest data, CMoE produces a well-designed, usable MoE from a 7B dense model within five minutes.
-With lightweight fine-tuning, it achieves high-performance recovery in under an hour.
+Scaling large language models (LLMs) improves performance but dramatically increases inference costs.
+The feed-forward network (FFN), consuming approximately 70% of inference compute, represents a critical bottleneck, particularly in large batch size scenarios.
+While mixture-of-experts (MoE) architectures leverage activation sparsity for efficiency, converting existing dense models to MoEs traditionally requires resource-intensive continual pre-training.
+We present CMoE, a framework that rapidly transforms dense LLMs into MoEs without training.
+The key innovation lies in analyzing FFN neuron activations to partition them into shared (always active) and routed experts.
+Routed neurons are clustered using a balanced assignment algorithm, and a differentiable router is constructed analytically from activation statistics, enabling immediate deployment or optional lightweight fine-tuning.
+Experiments demonstrate that, with activation ratio of 75%, it achieves remarkable results, delivering lossless precision in terms of perplexity while still maintaining a 5% acceleration.
+Further experiments reveal that a CMoE configuration activating just 25% of parameters reduces end-to-end latency by 1.5× while preserving usable perplexity without additional training.
+Moreover, a brief LoRA fine-tuning process (requiring only 1 hour and 2,000 samples) successfully recovers over 76% of the dense model’s downstream accuracy.
+By effectively balancing performance and efficiency, CMoE offers a viable path forward for deploying LLMs in real-world scenarios where computational resources are limited.
 We make our code publicly available at https://github.com/JarvisPei/CMoE.
-I Introduction
-Large language models (LLMs) have demonstrated exceptional proficiency in managing complex tasks and exhibiting emergent capabilities across diverse domains and applications, particularly when scaled to billions of parameters [1, 2, 3, 4].
-While LLMs have attained remarkable success, their expanding computational demands and model sizes have intensified challenges related to practical deployment, especially in environments with limited hardware resources or stringent latency requirements.
-To mitigate these challenges, mixture-of-experts (MoE) architectures [5, 6, 7, 8] have emerged as a promising paradigm.
-Unlike dense LLMs, where all parameters are activated for every input token, MoE models replace monolithic feed-forward networks (FFNs) with sparsely activated experts: specialized sub-networks that process inputs conditionally via dynamic routing.
-This design decouples model capacity from computational cost—activating only a subset of experts per token while preserving the model’s expressive power.
-Recently, researchers have found that there is high activation sparsity in the hidden neurons of FFNs of dense LLMs, which motivates them to develop sparsity-aware acceleration techniques to reduce computational overhead while maintaining model performance [9, 10].
-Building upon this insight, a growing body of research has focused on transforming dense LLMs into MoE architectures through strategic reorganization of FFN parameters but not training MoE from scratch [11, 12, 13].
-The prevailing methodology replaces conventional FFN layers with MoE layers, where neurons are partitioned into multiple expert sub-networks while maintaining the original parameter count.
-During inference, a routing mechanism selectively activates only a subset of experts per input token, thereby achieving dynamic sparsity without compromising model capacity.
-However, due to the high sparsity target always required in MoE models, these works often need massive computing resources and billions of training data for continual pre-training on the constructed MoE models.
-To address the aforementioned limitations, we propose carved MoE, named CMoE, a framework that efficiently carves sparse MoE architectures from dense LLMs through parameter reorganization and training-free structural adaptation.
-Unlike prior approaches that rebuild MoE models from scratch via resource-intensive pre-training, CMoE strategically “carves” experts from the dense model’s existing feed-forward network (FFN) neurons while preserving their inherent knowledge.
-This carving process is both computationally efficient—requiring only minutes of lightweight processing—and sophisticated, as it leverages systematic neuron grouping and analytical router construction to retain performance with minimal fine-tuning.
-The core innovation lies in CMoE’s ability to restructure dense FFNs into MoE layers without re-training the base model.
-First, we identify neurons that universally encode common knowledge (a.k.a.
-shared experts) and those that exhibit specialized, input-dependent activation patterns (a.k.a.
-routed experts).
-Shared experts are systematically retained by selecting neurons with the highest activation rates, ensuring they capture broadly applicable features.
-By formulating routed expert grouping as a balanced linear assignment problem, solved via the Jonker-Volgenant algorithm, CMoE clusters neurons into experts while maintaining parameter balance and activation coherence.
-Second, we derive the routing mechanism directly from the dense model’s activation statistics, bypassing the need for end-to-end router training.
-This involves constructing a differentiable routing function initialized using representative neurons from each expert cluster, enabling immediate usability while preserving optimization potential.
-In summary, the key contributions of this paper are:
-  • CMoE: A framework that efficiently carves MoE from dense LLMs by reorganizing FFN neurons into shared/routed experts, eliminating costly pre-training.
-  • Shared and routed experts: ensuring parameter balance and activation coherence for CMoE.
-  • Training-Free Routing: using representative neurons and lightweight adaptation for router initialization, which enables rapid performance recovery.
-Extensive experiments show that, with activation ratio $25\%$, CMoE can maintain a reasonable perplexity even without fine-tuning, and achieve $76.59\%$ accuracies of dense models on some downstream benchmarks with lightweight fine-tuning on 2,048 samples.
-II Related Work
+1 Introduction
+Large language models (LLMs) achieve remarkable performance zhang2022opt ; touvron2023llama ; liu2024visual ; liu2024deepseek , but their growing size presents deployment challenges due to high computational demands, especially on resource-constrained hardware or under strict latency budgets.
+The mixture-of-experts (MoE) architecture lepikhin2020gshard ; du2022glam ; fedus2022switch ; dai2024deepseekmoe offers a path forward.
+Instead of activating all parameters per token like dense models, MoEs use sparse feed-forward networks (FFNs) composed of specialized sub-networks (experts) selected by a dynamic router.
+This decouples model capacity from computational cost.
+In modern LLMs such as the Llama-series and DeepSeek-series, FFN blocks are responsible for approximately 70% of both inference FLOPs and memory traffic.
+As batch sizes increase, a common strategy for maximizing hardware utilization and throughput, models increasingly operate in a compute-bound regime, where computational power becomes the limiting factor for inference speed.
+Given the FFN’s disproportionate contribution to memory and computational demands, sparsifying FFNs has become a primary target for efficiency improvement.
+Several studies have reported high activation sparsity in FFN neurons dejavu ; moefication ; pei2024fusegpt , inspiring a wave of techniques that aim to restructure dense FFN layers into sparse mixtures of experts (MoEs) by reorganizing existing parameters llama-moe ; llama-moe-v2 ; zheng2024learn .
+These MoE conversions can dramatically reduce both memory usage and computational overhead, improving throughput, particularly in large-batch, latency-sensitive inference scenarios.
+However, despite their promise, such methods generally require substantial additional computation and large-scale datasets for continual pre-training or fine-tuning after conversion, in order to recover or maintain model quality.
+This retraining step can be prohibitively expensive and slow, posing a significant barrier to rapid deployment in production environments.
+As a result, there is growing interest in developing lightweight strategies that minimize retraining costs while delivering tangible inference speedups in real-world settings.
+To overcome this bottleneck, we propose CMoE, a framework to efficiently convert dense LLMs into MoE architectures without requiring any training.
+The conversion process leverages the existing knowledge within the dense model’s FFN and completes in under five minutes on a single GPU.
+CMoE first analyzes neuron activation patterns across a small calibration dataset to distinguish frequently active neurons (grouped into shared experts) from sparsely active ones (grouped into routed experts).
+Routed experts are formed by clustering neurons using a balanced assignment algorithm jonker1988shortest to maintain activation coherence.
+Crucially, CMoE then constructs a differentiable routing mechanism analytically from activation statistics, bypassing the need for router training and enabling immediate use or subsequent lightweight fine-tuning.
+Our contributions are:
+  • Training-free conversion.
+CMoE rapidly restructures dense FFNs into shared and routed experts using activation analysis and balanced assignment, requiring only minutes and no gradient updates.
+In practical implementation, it is found that with an activation ratio of 75%, it achieves remarkable results, delivering lossless precision in terms of perplexity while still maintaining a 5% acceleration.
+  • Analytical, differentiable router. The router is initialized directly from activation statistics, providing immediate functionality and a strong starting point for optional fine-tuning.
+  • Controllable performance-efficiency trade-off.
+CMoE allows adjusting the expert activation ratio.
+Experiments show a spectrum from 1.5×1.5$\times$1.5 × latency reduction (at 25% activation) to near-dense perplexity (at 75% activation).
+Minimal fine-tuning (1 hour, 2k samples) recovers >76% of dense model downstream accuracy.
+CMoE’s combination of speed, flexibility, and training-free nature presents a compelling solution for researchers and practitioners requiring efficient, low-overhead deployment of LLMs.
+2 Related Work
 In contrast to pretraining MoE models from scratch, recent research has investigated the feasibility of constructing MoE architectures by repurposing existing dense LLMs.
-Current methodologies for deriving MoE models from dense checkpoints generally follow two paradigms: (1) partitioning parameters of FFNs while preserving the original model’s total parameter count [14, 10], or (2) expanding the model’s overall capacity while retaining activation dimensions comparable to standard dense models [15, 16].
+Current methodologies for deriving MoE models from dense checkpoints generally follow two paradigms: (1) partitioning parameters of FFNs while preserving the original model’s total parameter count zuo2022moebert ; moefication ; yang2024xmoe , or (2) expanding the model’s overall capacity while retaining activation dimensions comparable to standard dense models komatsuzaki2022sparse ; wu2024parameter .
 This work prioritizes the former approach.
-Notably, MoEBERT [14] introduces an importance-driven strategy to transform FFNs into expert modules by strategically redistributing top-scoring neurons across specialized components.
-Concurrently, MoEfication [10] leverages the discovery of sparse activation patterns in ReLU-based FFNs within T5 architectures, enabling the decomposition of these layers into distinct expert groups governed by a learned routing mechanism.
-Based on continual training, the LLaMA-2 7B model is modified as a LLaMA-MoE-3.5B MoE model, where the parameters of the original FFNs are partitioned into multiple experts [11].
-After training with 200B tokens, the LLaMA-MoE-3.5B model significantly outperforms dense models that contain similar activation parameters.
-Furthermore, based on a two-stage post-training strategy, an MoE model is constructed from the LLaMA3 8B model, where both attention and MLP are partitioned into MoE blocks [12]].
-Extensive experiments have shown the effectiveness of constructing an MoE model from a dense model, and many techniques can be utilized to guarantee performance recovery.
-However, such performance recovery is extremely resource-consuming, which is unfavorable for efficient deployment in industrial applications.
-Therefore, more lightweight methods are required, such that performance recovery can be done within hours and even training-free.
-Note that model compression such as pruning and quantization is another important technique for efficient LLM inference [17, 18, 19, 20].
+Notably, MoEBERT zuo2022moebert introduces an importance-driven strategy to transform FFNs into expert modules by strategically redistributing top-scoring neurons across specialized components.
+Concurrently, MoEfication moefication leverages the discovery of sparse activation patterns in ReLU-based FFNs within T5 architectures, enabling the decomposition of these layers into distinct expert groups governed by a learned routing mechanism.
+Based on continual training, llama-moe modifies llama-moe modifies the LLaMA-2 7B model as a LLaMA-MoE-3.5B MoE model, where the parameters of the original FFNs are partitioned into multiple experts.
+In llama-moe-v2 , based on a two-stage post-training strategy, an MoE model is constructed from the LLaMA3 8B model, where both attention and MLP are partitioned into MoE blocks.
+Note that model compression such as pruning and quantization is another important technique for efficient LLM inference lin2024awq ; pei2023quantization ; zou2024bie ; lin2024qserve ; lu2024not .
 Pruning is among the most widely utilized approaches to detect and remove redundant or less significant parameters from models, thereby resulting in a sparser weight matrix and faster inference.
-ShortGPT [21] has put forward a simple layer-removal approach.
-This approach is based on block influence, which is determined by the similarity between a layer’s input and output.
-SliceGPT [22] substitutes each weight matrix with a smaller dense matrix, thereby decreasing the embedding dimension of the network.
-By taking into account contextual sparsity in real time, Deja Vu [9] and FuseGPT [23] have been proposed to accelerate LLM inference.
-In contrast to the post-training methods, Learn-To-be-Efficient is designed to train efficiency-aware LLMs so that they learn to activate fewer neurons and achieve a more favorable trade-off between sparsity and performance [24].
-III Background
-This study primarily focuses on the LLaMA family [2, 25], which uses SwiGLU [26] as the activation function.
-However, our analysis and findings can be adapted to most of the FFN structures of existing LLMs, including the ReLU-based FFNs [27].
-An FFN exists in the tail of each transformer block, which gets the input embedding $\mathbf{x}\in\mathbb{R}^{d}$ and then contributes to the output together with the residual connection, i.e.
+ShortGPT shortgpt determines the layer-removal by the similarity between a layer’s input and output.
+SliceGPT slicegpt substitutes each weight matrix with a smaller dense matrix oo deease the embedding dimension.
+By taking into account contextual sparsity in real time, Deja Vu dejavu has been proposed to accelerate LLM inference.
+In contrast to the post-training methods, Learn-To-be-Efficient is designed to train efficiency-aware LLMs so that they learn to activate fewer neurons and achieve a more favorable trade-off between sparsity and performance.
+3 Methodology
+Figure 1: The overview of our proposed CMoE.
+CMoE transforms dense LLMs into sparsely activated MoE architectures through two key phases: efficient expert grouping and training-free router construction, followed by optional lightweight adaptation.
+As shown in Figure 1(a), CMoE operates through the following systematic process:
+A.
+Neuron Activation Profiling (Section 3.1) Using a small calibration dataset, CMoE profiles the activation patterns of neurons within each FFN layer to categorize them into shared experts (high-activation, task-agnostic) and routed experts (sparsely activated, task-specific).
+B.
+Expert Grouping (Section 3.1) Shared Experts: Neurons exhibiting the highest activation rates are directly assigned to shared experts, which remain consistently activated during inference.
+Routed Experts: The remaining neurons are efficiently partitioned into routed experts through balanced clustering, mathematically formulated as a linear assignment problem.
+C.
+Router Construction (Section 3.2) The routing mechanism is analytically derived from the activation statistics of representative neurons in each expert cluster, effectively eliminating the need for computationally expensive end-to-end training.
+We further enhance CMoE by making the routing function differentiable (Section 3.3), facilitating further alignment and performance recovery by optional fine-tuning when resources permit.
+3.1 Shared and Routed Experts Grouping
+Note that an FFN exists in the tail of each transformer block, which gets the input embedding $\mathbf{x}\in\mathbb{R}^{d}$ and then contributes to the output together with the residual connection, i.e.
 $\mathbf{x}+F(\mathbf{x})$.
 Typically, an FFN is a two-layer fully connected network, i.e.
-the up projection and down projection layer, with an activation function between them.
+the up projection and down projection layers, with an activation function between them.
 For LLaMA, the SwiGLU composes another gate projection layer.
 Given the up projection weight $\mathbf{W}_{up}\in\mathbb{R}^{d\times d_{h}}$, the gate projection weight $\mathbf{W}_{gate}\in\mathbb{R}^{d\times d_{h}}$ and the down projection weight $\mathbf{W}_{down}\in\mathbb{R}^{d_{h}\times d}$, the process of an FFN is given by:
-$F(\mathbf{x})=\mathbf{h}\mathbf{W}_{down},$ (1)
-$\mathbf{h}\!=\!\text{SwiGLU}(\mathbf{x}\mathbf{W}_{up})\!=\!\text{Swish}(\mathbf{x}\mathbf{W}_{gate})\odot(\mathbf{x}\mathbf{W}_{up}),$
+$F(\mathbf{x})=\mathbf{h}\mathbf{W}_{down},\mbox{ and }\;\mathbf{h}=\text{SwiGLU}(\mathbf{x}\mathbf{W}_{up})=\text{Swish}(\mathbf{x}\mathbf{W}_{gate})\odot(\mathbf{x}\mathbf{W}_{up}),$ (1)
 where $\text{Swish}(x)=x\cdot\sigma(x)$ is element-wise and $\sigma(\cdot)$ is the sigmoid function.
-The basic MoE architecture is composed of a set of $N$ independent FFNs as experts, $\{E_{1},E_{2},...,E_{N}\}$, and a router network $G$ [5]. The output of an MoE-version FFN is then obtained by
-$\!\!\!\!F_{MoE}(\mathbf{x})$ $=\!\sum^{N}_{i=1}g_{i}E_{i}(\textbf{x}),$ (2)
-$g_{i}$ =$=\!\left\{\begin{array}[]{rl}s_{i},\!&s_{i}\in\text{TopK}(\{s_{i}|1\leq j\leq N\},K),\\ 0,\!&\text{otherwise},\end{array}\right.$
-s $=\!\left[s_{1},s_{2},\cdots,s_{N}\right]=G(\textbf{x}),$
-where $g_{i}$ is the score for the $i$-th expert, $\textbf{s}\in\mathbb{R}^{N}$ is the token-to-expert affinity, i.e.
-the output of $G$, and $\text{TopK}(\cdot,K)$ denotes the set comprising $K$ highest scores among the affinity scores calculated for $\mathbf{x}$ on all experts.
-IV Methodology
-[…]
-Figure 1: The overview of our proposed CMoE.
-IV-A Shared and Routed Experts Grouping
-CMoE starts with grouping neurons in FFN into experts.
+CMoE starts with grouping neurons in FFNs into experts.
+We construct $N$ experts of size $m=\frac{d_{h}}{N}$ (assuming the intermediate size $d_{h}$ is a multiple of $N$, i.e., $N\mid d_{h}$), including $N_{s}$ shared experts and $N_{r}$ routed experts $(N_{s}+N_{r}=N)$.
 Shared experts are expected to process common knowledge instead of specialization.
 Thus, the key idea is to group neurons that are always activated during FFN inference.
-For routed expert grouping, as described in previous works [28], the key idea is to cluster the neurons that are always activated simultaneously.
-CMoE carries forward these ideas but comes up with a detailed and analytical perspective.
-In this section, we construct $N$ experts of size $m=\frac{d_{h}}{N}$ (assuming $d_{h}$ is a multiple of $N$, i.e., $N\mid d_{h}$), including $N_{s}$ shared experts and $N_{r}$ routed experts $(N_{s}+N_{r}=N)$.
-We begin with analyzing the isolated contribution of each neuron to the output of FFN.
-Consider the $i$-th element of $\mathbf{h}\in\mathbb{R}^{d_{h}}$: $h_{i}=\text{Swish}(\mathbf{x}\cdot\mathbf{w}_{gate,i})\cdot(\mathbf{x}\cdot\mathbf{w}_{up,i})$, where $\mathbf{w}_{gate,i}\in\mathbb{R}^{d}$ and $\mathbf{w}_{up,i}\in\mathbb{R}^{d}$ are the $i$-th column of $\mathbf{W}_{gate}$ and $\mathbf{W}_{up}$, respectively.
+Given $\mathbf{x}\in\mathbb{R}^{d}$ the input embedding, consider the $i$-th element of hidden state $\mathbf{h}\in\mathbb{R}^{d_{h}}$: $h_{i}=\text{Swish}(\mathbf{x}\cdot\mathbf{w}_{gate,i})\cdot(\mathbf{x}\cdot\mathbf{w}_{up,i})$, where $\mathbf{w}_{gate,i}\in\mathbb{R}^{d}$ and $\mathbf{w}_{up,i}\in\mathbb{R}^{d}$ are the $i$-th column of $\mathbf{W}_{gate}\in\mathbb{R}^{d\times d_{h}}$ (gate projection weight) and $\mathbf{W}_{up}\in\mathbb{R}^{d\times d_{h}}$ (up projection weight), respectively.
 It shows that the value of $h_{i}$ only depends on the input $\mathbf{x}$ and the corresponding $i$-th weights, which implies the independence between neurons.
@@
 $F(\mathbf{x})=\sum_{i=1}^{d_{h}}h_{i}\mathbf{w}_{down,i},$ (2)
-where $\mathbf{w}_{down,i}\in\mathbb{R}^{d}$ is the $i$-th row of $\mathbf{W}_{down}$.
+where $\mathbf{w}_{down,i}\in\mathbb{R}^{d}$ is the $i$-th row of $\mathbf{W}_{down}\in\mathbb{R}^{d_{h}\times d}$ (down projection weight).
 When we revisit the FFN process, we can regard each $h_{i}$ as the score to be multiplied to the split vector $\mathbf{w}_{down,i}$, whose product contributes to part of the output $F(\mathbf{x})$.
-As a finding of some structured pruning research [29, 30], the norm of $F(\mathbf{x})$ is always small due to the residual connection.
-Such phenomenon implies the high sparsity existed in FFN, and by (3) we relate it to $h_{i}$, whose value decides how large an isolated neuron contributes to the output.
-Therefore we make a hypothesis:
+Therefore, we make a hypothesis:
 $\arg\min_{i}|h_{i}\mathbf{w}_{down,i}|\approx\arg\min_{i}|h_{i}|,$ (3)
 which is reasonable since when $h_{i}$ is extremely small, the product $h_{i}\mathbf{w}_{down,i}$ will also vanish.
-It is expected that the hidden state $\mathbf{h}$ should be highly sparse, which means $|h_{i}|$ is often extremely small.
-To verify it, we hack into the FFN and draw the distribution of $\mathbf{h}$.
-Figure 2: The histogram of FFN hidden state $\mathbf{h}$ for the $3$-th block and the $1,000$-th token.
-As demonstrated in fig.
-2, the distribution is sharply peaked at 00 and constrained within a small range, indicating that most $h_{i}$ are concentrated near zero and confirming the sparsity.
-And it exhibits symmetry, suggesting a near-normal distribution centered around the mean.
-Based on what we discuss and observe above, the hidden state values work well as the basis for judgment of neuron activation, because of its high differentiation and independence across different neurons.
+In Appendix A, we conduct experiments to demonstrate that the hidden state $\mathbf{h}$ is usually highly sparse, which means $|h_{i}|$ is often extremely small.
 Therefore, we propose a new metric, called absolute TopK (ATopK), to determine the activation status of a neuron with index $i$:
@@
 where we choose neurons with $K_{a}$ highest absolute value among the hidden state values, and assign their labels in the activation marker $\mathbf{a}=\left[a_{1},a_{2},\cdots,a_{d_{h}}\right]$ with $1$.
-To further evaluate the activation information, we make samples in the training set to record their activation markers.
-Given a batched input tensor $\mathbf{X}\in\mathbb{R}^{b\times s\times d}$, with $b$ the batch size and $s$ the sequence length, we obtain the batched hidden state $\mathbf{H}\in\mathbb{R}^{b\times s\times d_{h}}$ as follows:
-$\mathbf{H}=\text{Swish}(\mathbf{X}\mathbf{W}_{gate})\odot(\mathbf{X}\mathbf{W}_{up}).$ (6)
-Note that in practical implementation, we normalize $\mathbf{X}$, $\mathbf{W}_{gate}$ and $\mathbf{W}_{up}$ before the calculation to eliminate the influence of their magnitudes on the output.
-We then calculate the activation markers for the hidden state of all the tokens, which is reshaped as an activation feature matrix $\mathbf{A}\in\mathbb{R}^{q\times d_{h}},q=b\cdot s$.
-Denote the $i$-th column of $\mathbf{A}$ as the feature vector $\mathbf{c}_{i}\in\mathbb{R}^{q}$, i.e., $\mathbf{A}=[\mathbf{c}_{1}\;\mathbf{c}_{2}\;\cdots\;\mathbf{c}_{d_{h}}]$ which represents the $i$-th neuron’s activation status on the sampled $q$ tokens.
-By calculating the expected value of each $\mathbf{c}_{i}$, we can obtain the vector of activated rates $\bm{\mu}$ as:
-$\bm{\mu}=[\mu_{1},\mu_{2},\cdots,\mu_{d_{h}}],\mbox{ with }\mu_{i}=\frac{1}{q}\sum_{j=1}^{q}c_{j,i},$ (7)
-where $c_{j,i}$ indicates the $j$-th element of $\mathbf{c}_{i}$. We draw the histogram of $\bm{\mu}$ as in Fig. 3.
-Figure 3: The histogram of activation rates $\bm{\mu}$ for the $3$-th block with $K_{a}=1,000$.
-The histogram reveals a highly skewed distribution of activation rates, where the majority of neurons exhibit low activation rates (below 0.1), with a sharp peak near 0.07.
-However, the distribution also features a long tail, indicating the presence of a subset of neurons with significantly higher activation rates extending up to 1.
-These high-activation neurons are likely active across a wide range of input tokens, making them suitable for processing common knowledge rather than task-specific specialization.
-Therefore, we construct shared experts by grouping these high-activation neurons.
-Given the total number of shared experts as $N_{s}$ and the expert size $m$, we get the selection indices set $S_{N_{s}}$ by selecting $N_{s}\cdot m$ neurons with highest activation rates:
-$S_{N_{s}}=\{i\ :\ \mu_{i}\in\text{TopK}(\{\mu_{j}\mid 1\leq j\leq d_{h}\},N_{s}\cdot m)\}.$ (8)
-Since the independency across neurons, as we discussed, we build the experts using the same architecture as the original FFN and assign the expert parameters to the neurons of selected indices.
-Since all shared experts are always activated, we just construct them together for convenience:
-$\mathbf{W}_{up}^{s}$ $=\mathbf{W}_{up}[:,S_{N_{s}}],$ (9)
-$\mathbf{W}_{gate}^{s}$ $=\mathbf{W}_{gate}[:,S_{N_{s}}],$
-$\mathbf{W}_{down}^{s}$ $=\mathbf{W}_{down}[S_{N_{s}},:],$
-where $\mathbf{W}_{up}^{s}\in\mathbb{R}^{d\times N_{s}\cdot m}$, $\mathbf{W}_{gate}^{s}\in\mathbb{R}^{d\times N_{s}\cdot m}$ and $\mathbf{W}_{down}^{s}\in\mathbb{R}^{N_{s}\cdot m\times d}$ are the weights of linear layers of the shared experts $E^{s}$.
-Given the input embedding $\mathbf{x}$, the output of $E^{s}$ is obtained via
-$E^{s}(\mathbf{x})\!=\!\mathbf{h}^{s}\mathbf{W}_{down}^{s},\mbox{with }\mathbf{h}^{s}\!=\!\text{Swish}(\mathbf{x}\mathbf{W}_{gate}^{s})\!\odot\!(\mathbf{x}\mathbf{W}_{up}^{s}).$
-The majority of low activation rates also encourage us to construct routed experts, which are not always activated but are specialized for tokens encountered.
-To construct $N_{r}$ routed experts, we develop a customized balanced K-means clustering algorithm.
-We first identify $N_{r}$ centroids as the neurons (excluding those assigned to shared experts) with the highest activation rates, and group their feature vectors into a centroid set $C=\{\mathbf{c}_{i}:\mu_{i}\in\text{TopK}(\{\mu_{j}\mid 1\leq j\leq d_{h},j\notin S_{Ns}\},N_{r})\}=\{\hat{\mathbf{c}}_{1},\dots,\hat{\mathbf{c}}_{N_{r}}\}$, where, for convenience, we re-label the centroids with $\hat{\mathbf{c}}$.
-By selecting these centroids, we ensure that the clusters are initialized around meaningful and prominent activation patterns, providing a strong starting point for clustering.
-Using the centroids selected, we will construct clusters with remaining neurons.
-We pay attention to their correlation to the centroids, i.e.
-we study the distances of their feature vectors to the centroids’.
-The feature vector $\mathbf{c}$ individually contains the activation status of a specific neuron, and we expect that neurons with similar functionalities will have similar patterns on the activation status, i.e.
-they are often activated and deactivated together during inference.
-Therefore, we compute the $L_{2}$ distance of each feature vector $\mathbf{c}$ to each centroid $\hat{\mathbf{c}}$.
-Define a distance matrix $\mathbf{D}\in\mathbb{R}^{N_{r}\cdot m\times N_{r}}$, whose element $d_{i,j}$ is the $L_{2}$ distance between the $i$-th feature vector $\mathbf{c}_{i}$ and the $j$-th centroid $\hat{\mathbf{c}}_{j}$:
-$d_{i,j}=\|\mathbf{c}_{i}-\hat{\mathbf{c}}_{j}\|_{2}=\sqrt{\sum_{k=1}^{q}(c_{k,i}-\hat{c}_{k,j})^{2}},$ (10)
-where, for convenience, we also include the original feature vectors of centroids inside the set to be grouped.
-To group the neurons into routed experts, with the centroids designed above at $t=0$, the constrained balanced $K$-means clustering algorithm works as follows [31].
-Given $N_{r}\cdot m$ vectors $\mathbf{c}$, the cluster membership value $m$ and cluster centroids $\hat{\mathbf{c}}_{1}^{t},\hat{\mathbf{c}}_{2}^{t},\dots,\hat{\mathbf{c}}_{N_{r}}^{t}$ at iteration $t$, compute $\hat{\mathbf{c}}_{1}^{t+1},\hat{\mathbf{c}}_{2}^{t+1},\dots,\hat{\mathbf{c}}_{N_{r}}^{t+1}$ at iteration $t+1$ with the following $2$ steps:
-Cluster Assignment Let $T_{i,p}^{t}$ be a solution to the following linear program with $\hat{\mathbf{c}}_{p}^{t}$ fixed:
-$\min_{T}$ $\sum_{i=1}^{N_{r}\cdot m}\sum_{p=1}^{N_{r}}T_{i,p}\cdot d_{i,p}$ (11)
-s.t. $\sum_{i=1}^{N_{r}\cdot m}T_{i,p}=m,p=1,\dots N_{r}$
-$\sum_{p=1}^{N_{r}}T_{i,p},i=1,\dots,N_{r}\cdot m$
-$T_{i,p}\geq 0,i=1,\dots,N_{r}\cdot m,p=1,\dots N_{r}.$
-Cluster Update Update $\hat{\mathbf{c}}_{p}^{t+1}$ as follows:
+We then obtain the set of selection indices $S_{N_{s}}$ by selecting $N_{s}\cdot m$ neurons with the highest activation rates.
+And we construct the shared experts as follows:
+$\mathbf{W}_{up}^{s}=\mathbf{W}_{up}[:,S_{N_{s}}];\;\;\mathbf{W}_{gate}^{s}=\mathbf{W}_{gate}[:,S_{N_{s}}];\;\;\mathbf{W}_{down}^{s}=\mathbf{W}_{down}[S_{N_{s}},:];$ (5)
+$\Longrightarrow E^{s}(\mathbf{x})=\mathbf{h}^{s}\mathbf{W}_{down}^{s},\mbox{ with }\;\mathbf{h}^{s}=\text{Swish}(\mathbf{x}\mathbf{W}_{gate}^{s})\!\odot\!(\mathbf{x}\mathbf{W}_{up}^{s}).$
+Here, $\mathbf{W}_{up}^{s}\in\mathbb{R}^{d\times N_{s}\cdot m}$, $\mathbf{W}_{gate}^{s}\in\mathbb{R}^{d\times N_{s}\cdot m}$ and $\mathbf{W}_{down}^{s}\in\mathbb{R}^{N_{s}\cdot m\times d}$ are the weights of linear layers of the shared experts $E^{s}$.
+More detailed analysis is given in Appendix A.
+To construct $N_{r}$ routed experts, the key idea is to cluster the neurons that are always activated simultaneously.
+We first identify $N_{r}$ centroids by selecting neurons (excluding those already assigned to shared experts) with the highest activation rates.
+We group their feature vectors into a centroid set $C=\{\mathbf{c}_{i}:\mu_{i}\in\text{TopK}({\mu_{j}\mid 1\leq j\leq d_{h},j\notin S_{N_{s}}},N_{r})\}=\{\hat{\mathbf{c}}_{1},\dots,\hat{\mathbf{c}}_{N_{r}}\}$, where $\boldsymbol{\mu}=[\mu_{1},\mu_{2},\cdots,\mu_{d_{h}}]$ is the activation rate averaged from the marker $\mathbf{a}$ and we relabel the centroids as $\hat{\mathbf{c}}$.
+Each feature vector $\mathbf{c}$ encapsulates the activation status of a specific neuron across the calibration dataset.
+We hypothesize that neurons with similar functionalities exhibit similar activation patterns—specifically, they tend to be activated and deactivated together during inference.
+Based on this principle, we compute the $L_{2}$ distance between each feature vector $\mathbf{c}$ and each centroid $\hat{\mathbf{c}}$.
+We formalize this by constructing a distance matrix $\mathbf{D}\in\mathbb{R}^{N_{r}\cdot m\times N_{r}}$, where element $d_{i,j}$ represents the $L_{2}$ distance between the $i$-th feature vector $\mathbf{c}_{i}$ and the $j$-th centroid $\hat{\mathbf{c}}_{j}$:
+$d_{i,j}=\|\mathbf{c}_{i}-\hat{\mathbf{c}}_{j}\|_{2}=\sqrt{\sum_{k=1}^{q}(c_{k,i}-\hat{c}_{k,j})^{2}}.$ (6)
+To group the neurons into routed experts, with the centroids designed above at $t=0$, we adopt the constrained balanced $K$-means clustering method.
+Given $mN_{r}$ vectors $\mathbf{c}$, the cluster membership value $m$ and cluster centroids $\hat{\mathbf{c}}_{1}^{t},\hat{\mathbf{c}}_{2}^{t},\dots,\hat{\mathbf{c}}_{N_{r}}^{t}$ at iteration $t$, compute $\hat{\mathbf{c}}_{1}^{t+1},\hat{\mathbf{c}}_{2}^{t+1},\dots,\hat{\mathbf{c}}_{N_{r}}^{t+1}$ at iteration $t+1$ with the following $2$ steps:
+Cluster Assignment. Let $T_{i,p}^{t}$ be a solution to the following linear program with $\hat{\mathbf{c}}_{p}^{t}$ fixed:
+$\min_{T}$ $\sum_{i=1}^{mN_{r}}\sum_{p=1}^{N_{r}}T_{i,p}\cdot d_{i,p}$ (7)
+s.t. $\sum_{i=1}^{mN_{r}}T_{i,p}=m,\forall p\in\{1,\dots,N_{r}\};\;\sum_{p=1}^{N_{r}}T_{i,p}=1,\;\forall i\in\{1,\dots,mN_{r}\};\;T_{i,p}\geq 0,\forall p,i.$
+Cluster Update. Update $\hat{\mathbf{c}}_{p}^{t+1}$ as follows:
 $\hat{\mathbf{c}}_{p}^{t+1}=\left\{\begin{array}[]{ll}\frac{\sum_{i=1}^{N_{r}\cdot m}T_{i,p}^{t}\cdot\mathbf{c}_{i}}{\sum_{i=1}^{N_{r}\cdot m}T_{i,p}^{t}},&\text{if}\sum_{i=1}^{N_{r}\cdot m}T_{i,p}^{t}>0,\\ \hat{\mathbf{c}}_{p}^{t},&\text{otherwise}.\end{array}\right.$ (8)
-The steps are terminated until $\hat{\mathbf{c}}_{p}^{t+1}=\hat{\mathbf{c}}_{p}^{t}$, $\forall p=1,\dots,N_{r}$.
+The stopping criterion is $\hat{\mathbf{c}}_{p}^{t+1}=\hat{\mathbf{c}}_{p}^{t}$, $\forall p=1,\dots,N_{r}$.
 We include the distance value $d_{i,j}$ into the cost function so that the balanced $K$-means clustering makes the intra-expert distances low and inter-expert distances high.
-However, the cluster assignment we defined above is an unbalanced assignment problem ($N_{r}\cdot m>m$) and cannot directly be solved by existing algorithms for balanced assignment.
+However, the cluster assignment defined above is an unbalanced assignment problem ($mN_{r}>m$) and cannot be solved directly by existing algorithms for balanced assignment.
 Therefore, we reduce it to a balanced assignment by extending the distance matrix $\mathbf{D}$ as follows:
@@
 Then the reduced balanced assignment problem is formulated as:
-Balanced Assignment Let $T_{i,p^{\prime}}^{{}^{\prime},t}$ be a solution to the following linear program with $\hat{\mathbf{c}}_{p}^{t}$ fixed:
-$\min_{T^{\prime}}$ $\sum_{i=1}^{N_{r}\cdot m}\sum_{p^{\prime}=1}^{N_{r}\cdot m}T^{\prime}_{i,p^{\prime}}\cdot d_{i,p^{\prime}}^{ext}$ (13)
-s.t. $\sum_{i=1}^{N_{r}\cdot m}T^{\prime}_{i,p^{\prime}}=1,p^{\prime}=1,\dots N_{r}\cdot m$
-$\sum_{p^{\prime}=1}^{N_{r}\cdot m}T^{\prime}_{i,p^{\prime}}=1,i=1,\dots,N_{r}\cdot m$
-$T^{\prime}_{i,p^{\prime}}\geq 0,i=1,\dots,N_{r}\cdot m,p^{\prime}=1,\dots N_{r}\cdot m,$
-where $d_{i,p^{\prime}}^{ext}=d_{i,p}$ if $p\cdot m<p^{\prime}\leq(p+1)\cdot m$.
-And the balanced cluster $\hat{\mathbf{c}}_{p}^{t+1}$ can be updated as follows:
+Balanced Assignment. Let $T_{i,p^{\prime}}^{{}^{\prime},t}$ be a solution to the following linear program with $\hat{\mathbf{c}}_{p}^{t}$ fixed:
+$\min_{T^{\prime}}$ $\sum_{i=1}^{mN_{r}}\sum_{p^{\prime}=1}^{N_{r}\cdot m}T^{\prime}_{i,p^{\prime}}\cdot d_{i,p^{\prime}}^{ext}$ (9)
+s.t.
+$\sum_{i=1}^{mN_{r}}T^{\prime}_{i,p^{\prime}}=1,\forall p^{\prime}\in\{1,\dots,mN_{r}\};\;\sum_{p^{\prime}=1}^{mN_{r}}T^{\prime}_{i,p^{\prime}}=1,\forall i\in\{1,\dots,mN_{r}\};\;T^{\prime}_{i,p^{\prime}}\geq 0,\forall I,p^{\prime},$
+where $d_{i,p^{\prime}}^{ext}=d_{i,p}$ if $mp<p^{\prime}\leq m(p+1)$. The balanced cluster $\hat{\mathbf{c}}_{p}^{t+1}$ can be updated as follows:
 $\hat{\mathbf{c}}_{p}^{t+1}=\scriptstyle\left\{\begin{array}[]{ll}\frac{\sum\limits_{k=k_{0}}^{mp}\sum\limits_{i=1}^{N_{r}\cdot m}T_{i,k}^{{}^{\prime},t}\cdot\mathbf{c}_{i}}{\sum\limits_{k=k_{0}}^{mp}\sum\limits_{i=1}^{N_{r}\cdot m}T_{i,k}^{{}^{\prime},t}},&\text{if}\scriptstyle\sum\limits_{k=k_{0}}^{mp}\sum\limits_{i=1}^{N_{r}\cdot m}T_{i,k}^{{}^{\prime},t}>0,\\ \hat{\mathbf{c}}_{p}^{t},&\text{otherwise},\end{array}\right.$ (10)
 where $k_{0}=m(p-1)+1$.
-Drawing on the Jonker-Volgenant algorithm [32], this problem can be addressed as a reduced assignment problem in each step of the $K$-means algorithm, with a complexity of $O(n^{3})$.
-The final solution $T^{\prime}$ gives us the optimized strategy to group the routed experts. We get the selection indices set $S_{N_{r},p}$, $p=1,\dots,N_{r}$, for each routed expert $E^{r}_{p}$:
+The final solution $T^{\prime}$ gives us the optimized strategy to group the routed experts.
+We get the selection indices set $S_{N_{r},p}$, $p=1,\dots,N_{r}$, for each routed expert $E^{r}_{p}$:
 $S_{N_{r},p}=\{i\ :\exists\ T^{\prime}_{i,k}=1,\ \text{for}\ \forall k\in\{m(p-1)+1,\dots,m\}\}.$
-We then build the weights $\mathbf{W}_{up}^{r,p}\in\mathbb{R}^{d\times m}$, $\mathbf{W}_{gate}^{r,p}\in\mathbb{R}^{d\times m}$ and $\mathbf{W}_{down}^{r,p}\in\mathbb{R}^{m\times d}$ of each routed expert $E_{p}^{r}$ as in Equation 9.
+We then build the weights $\mathbf{W}_{up}^{r,p}\in\mathbb{R}^{d\times m}$, $\mathbf{W}_{gate}^{r,p}\in\mathbb{R}^{d\times m}$ and $\mathbf{W}_{down}^{r,p}\in\mathbb{R}^{m\times d}$ of each routed expert $E_{p}^{r}$ as in (LABEL:eq:expert_w).
 And the output of $E_{p}^{r}$ is given by:
@@
 We expect the router to choose the routed experts with top $N_{k}$ scores.
-Therefore, we modify the MoE-version FFN as in (IV-A),
+Therefore, we modify the MoE-version FFN as in (3.1),
 $F_{MoE}(\mathbf{x})=E^{s}(\mathbf{x})+\sum^{N_{r}}_{i=1}g_{i}E_{i}^{r}(\textbf{x}),$
-$g_{i}$ =$=\left\{\begin{array}[]{rl}1,&s_{i}\in\text{TopK}(\{s_{i}|1\leq j\leq N_{r}\},N_{k}),\\ 0,&\text{otherwise},\end{array}\right.$ (17)
-s $=\left[s_{1},s_{2},\cdots,s_{N_{r}}\right]=G(\textbf{x}),$ (18)
+$g_{i}=\left\{\begin{array}[]{rl}1,&s_{i}\in\text{TopK}(\{s_{i}|1\leq j\leq N_{r}\},N_{k}),\\ 0,&\text{otherwise},\end{array}\right.\mbox{ and }\;\textbf{s}=\left[s_{1},s_{2},\cdots,s_{N_{r}}\right]=G(\textbf{x}),$ (13)
 where we make the expert score $g_{i}\in\{0,1\}$ to enable an un-scaled version of the expert output to avoid biases.
-IV-B Training-free Router Construction
-We now present the training-free router network $G$ for CMoE.
+3.2 Training-free Router Construction
 Unlike previous works, which either built the router from scratch or intuitively used hidden features as initialization, we formulate the router construction as a minimization problem and then develop an algorithm with analysis to construct the router by approximating the optimal solution.
-Given the same input embedding $\mathbf{x}$, the output of original output of the dense FFN, i.e. $F(\mathbf{x})$ in (1), is equivalent to the sum of the output of all the experts in $F_{MoE}$:
-$F(\mathbf{x})$ $=E^{s}(\mathbf{x})+\sum^{N_{r}}_{i=1}E_{i}^{r}(\textbf{x}).$ (19)
-The only difference between (19) and $F_{MoE}(\mathbf{x})$ in (IV-A) is the expert score $\mathbf{g}$, which is obtained from the TopK selection of the output of $G$.
-Therefore, to preserve important knowledge captured by the original dense FNN, $G$ can be constructed to enforce $F_{MoE}(\mathbf{x})$ to be close to $F(\mathbf{x})$ by solving the minimization problem (IV-B),
-$\arg\min_{G}|F_{MoE}(x;G)-F(x)|$
-=\!= $\!\arg\min_{G}|\sum^{N_{r}}_{i=1}(g_{i}-1)E_{i}^{r}(\textbf{x})|=\!\!\arg\min_{G}|\sum^{i\in S_{de}}E_{i}^{r}(\mathbf{x})|,$
+Given the same input embedding $\mathbf{x}$, the output of original output of the dense FFN is equivalent to the sum of the output of all the experts in $F_{MoE}$: $F(\mathbf{x})=E^{s}(\mathbf{x})+\sum^{N_{r}}_{i=1}E_{i}^{r}(\textbf{x})$.
+The only difference roots in the expert score $\mathbf{g}$, which is obtained from the TopK selection of the output of $G$.
+Therefore, to preserve important knowledge captured by the original dense FNN, $G$ can be constructed to enforce $F_{MoE}(\mathbf{x})$ to be close to $F(\mathbf{x})$ by solving the minimization problem (14),
+$\arg\min_{G}|F_{MoE}(x;G)-F(x)|=\arg\min_{G}|\sum^{N_{r}}_{i=1}(g_{i}-1)E_{i}^{r}(\textbf{x})|=\arg\min_{G}|\sum^{i\in S_{de}}E_{i}^{r}(\mathbf{x})|,$ (14)
 where $S_{de}\!=\!\{i:s_{i}\!\notin\!\text{TopK}(\{s_{i}|1\!\leq\!j\!\leq\!N_{r}\},N_{k})\}$ and $|S_{de}|\!=\!N_{r}\!-\!N_{k}$, and the problem becomes constructing the $G$ to minimize the absolute sum of the output of deactivated routed experts.
 Note that we have made a hypothesis in (3) that the output/sparsity of $F(x)$ is highly related to the norm/sparsity of $\mathbf{h}$, which is the same for the expert outputs.
-Based on (3) and (4), we reformulate (IV-B) as in (20):
-$\arg\min_{G}|\sum^{i\in S_{de}}E_{i}^{r}(\mathbf{x})|$ (20)
+We reformulate (14) as in (15):
 […]
 […]
-=\!= $\!\arg\min_{G}\mathbb{E}_{\mathbf{h}}\left[\|\mathbf{h}^{r}_{i}\|_{1}\mid i\in S_{de}\right].$
 The problem becomes constructing $G$ that can minimize the expected hidden states $\mathbf{h}$ of deactivated routed experts.
@@
 $\exists\ \text{permutation}\ \sigma$ such that:
-$s_{\sigma(1)}\leq s_{\sigma(2)}\leq\dots\leq s_{\sigma(N_{r})}\mbox{ and}$
-$\bar{\mathbf{h}}^{r}_{\sigma(1)}\leq\bar{\mathbf{h}}^{r}_{\sigma(2)}\leq\dots\leq\bar{\mathbf{h}}^{r}_{\sigma(N_{r})},$ (21)
+$s_{\sigma(1)}\leq s_{\sigma(2)}\leq\dots\leq s_{\sigma(N_{r})}\mbox{ and}\ \ \ \bar{\mathbf{h}}^{r}_{\sigma(1)}\leq\bar{\mathbf{h}}^{r}_{\sigma(2)}\leq\dots\leq\bar{\mathbf{h}}^{r}_{\sigma(N_{r})},$ (16)
 by which we can verify that the minimum of (15) is:
@@
 which is obtained by setting $G(\mathbf{x})$ as:
-$S_{de}$ =$=\{\sigma(1),\dots,\sigma(N_{r}-N_{k})\}$
-=$=\{i:s_{i}\notin\text{TopK}(\{s_{i}|1\leq j\leq N_{r}\},N_{k})\}.$ (23)
-Finally, we pay attention to the hidden state $\mathbf{h}^{r}$ of any routed expert, which is the output by the neurons that we grouped in Section IV-A, where the cluster is centered to the centroid $\hat{\mathbf{c}}$.
-We denote the neuron in each cluster that has the lowest $L_{2}$ distance to the centroid as the representative neuron:
+$S_{de}=\{\sigma(1),\dots,\sigma(N_{r}-N_{k})\}=\{i:s_{i}\notin\text{TopK}(\{s_{i}|1\leq j\leq N_{r}\},N_{k})\}.$ (18)
+The neuron closest to the centroid in each cluster is considered as the representative neuron:
 $R_{j}=i,\mbox{ if }\|\mathbf{c}_{i}-\hat{\mathbf{c}}_{j}\|_{2}\leq\|\mathbf{c}_{k}-\hat{\mathbf{c}}_{j}\|_{2},\forall k\in S_{N_{r},j}.$ (19)
@@
 where $\mathbf{W}_{gate}^{R}=\mathbf{W}_{gate}[:,S_{R}]$, $\mathbf{W}_{up}^{R}=\mathbf{W}_{up}[:,S_{R}]$, and $S_{R}=\{R_{1},\dots,R_{N_{r}}\}$. This leads to
-$G(\mathbf{x})$ $=\left[s_{1},s_{2},\cdots,s_{N_{r}}\right]=\left[h^{r}_{R_{1}},h^{r}_{R_{2}},\cdots,h^{r}_{R_{N_{r}}}\right]$
-$\approx\left[\bar{\mathbf{h}}^{r}_{1},\mathbf{h}^{r}_{2},\cdots,\mathbf{h}^{r}_{N_{r}}\right],$ (26)
-which is hence an approximate solution for the original problem introduced in (IV-B).
-IV-C Differentiable Routing and Load-balancing
-Though we have constructed a well-designed router in Section IV-B, it is not differentiable since each expert score $g_{i}$ is a constant, as shown in (IV-A), hindering further alignment and performance recovery.
+$G(\mathbf{x})=\left[s_{1},s_{2},\cdots,s_{N_{r}}\right]=\left[h^{r}_{R_{1}},h^{r}_{R_{2}},\cdots,h^{r}_{R_{N_{r}}}\right]\approx\left[\bar{\mathbf{h}}^{r}_{1},\mathbf{h}^{r}_{2},\cdots,\mathbf{h}^{r}_{N_{r}}\right],$ (21)
+which is hence an approximate solution for the original problem introduced in (14).
+3.3 Differentiable Routing and Load-balancing
+Though we have constructed a well-designed router in Section 3.2, it is not differentiable since each expert score $g_{i}$ is a constant, as shown in (3.1), hindering further alignment and performance recovery.
 Therefore, we introduce a learnable parameter $\mathbf{u}$ when computing the expert scores as follows,
-$g_{i}\!$ =$\!=\!\left\{\begin{array}[]{ll}\!\!1+s^{\prime}_{i}\cdot u_{i},&s^{\prime}_{i}\in\text{TopK}(\{s^{\prime}_{i}|1\leq j\leq N_{r}\},N_{k}),\\ \!\!0,&\text{otherwise},\end{array}\right.$ (28)
-$\mathbf{s}^{\prime}$ $=\text{Softmax}(\mathbf{s}),\ \textbf{u}=[u_{1},u_{2},\dots,u_{N_{r}}]$ (29)
+$g_{i}\!\!=\!\left\{\begin{array}[]{ll}\!\!1+s^{\prime}_{i}\cdot u_{i},&s^{\prime}_{i}\in\text{TopK}(\{s^{\prime}_{i}|1\leq j\leq N_{r}\},N_{k}),\\ \!\!0,&\text{otherwise},\end{array}\right.\mathbf{s}^{\prime}=\text{Softmax}(\mathbf{s}),\;\&\;\textbf{u}=[u_{1},u_{2},\dots,u_{N_{r}}],$ (23)
 where the scale $\mathbf{u}$ is initialized as zero to avoid perturbation.
-For MoE models, load-balancing is crucial to guarantee computational efficiency, especially expert parallelism in LLMs serving.
-As in DeepSeek-V3 [4], we use the auxiliary-loss-free load balancing by introducing a bias term $\mathbf{b}$ before the TopK selection:
-$g_{i}\!$ =$\!=\!\left\{\begin{array}[]{ll}\!\!1+s^{\prime}_{i}\cdot u_{i},&s^{\prime}_{i}+b_{i}\in\text{TopK}(\{s^{\prime}_{i}|1\leq j\leq N_{r}\},N_{k}),\\ \!\!0,&\text{otherwise},\end{array}\right.$ (32)
-$\mathbf{b}$ $=[b_{1},b_{2},\dots,b_{N_{r}}].$ (33)
+For load balancing, we introduce a bias term $\mathbf{b}$ before the TopK selection as follows:
+$g_{i}\!\!=\!\left\{\begin{array}[]{ll}\!\!1+s^{\prime}_{i}\cdot u_{i},&s^{\prime}_{i}+b_{i}\in\text{TopK}(\{s^{\prime}_{i}|1\leq j\leq N_{r}\},N_{k}),\\ \!\!0,&\text{otherwise},\end{array}\right.\mbox{ and }\;\mathbf{b}=[b_{1},b_{2},\dots,b_{N_{r}}].$ (26)
 Here, $\mathbf{b}$ is initialized as zero and updated based on the expert load status during each step of training, as in DeepSeek-V3.
@@
 decreasing/increasing the bias term by $\gamma$ if the corresponding expert is overloaded/underloaded.
-V Experiments
-TABLE I: Comparison of perplexity results. Randomly select samples from WikiText-2 training dataset for calibration and training.
-Method Type LLaMA-2-7B LLaMA-3-8B Training-free Fine-tuning Training-free Fine-tuning WikiText-2 C4 WikiText-2 C4 WikiText-2 C4 WikiText-2 C4 Dense - 5.27 7.27 - - 6.14 9.44 - - LLaMA-MoE A2E8 nan nan 468.00 2660.68 nan nan 988.20 7521.83 LLaMA-MoE A4E16 nan nan 540.62 2690.68 nan nan 1094.24 7758.91 CMoE S1A1E8 60.86 135.61 12.76 32.12 162.74 324.71 21.16 64.03 CMoE S1A3E16 89.19 180.65 13.84 33.56 262.85 465.09 22.97 72.34 CMoE S2A2E16 62.30 136.12 12.73 32.37 143.38 284.19 21.01 65.57
-CMoE is implemented based on Hugging Face Transformers [33] together with Pytorch [34].
-The experiments are conducted on one NVIDIA H800 PCIe 80GB graphics card with CUDA Driver 12.6.
-We randomly select samples from WikiText-2 training dataset [35] as calibration and fine-tuning data.
-We use only 8 examples with 2,048 sequence length as calibration for calculating the hidden state $\mathbf{H}$ in (6).
+4 Experiments
+CMoE is implemented using Hugging Face Transformers wolf2019huggingface and PyTorch paszke2019pytorch .
+We use samples from WikiText-2 merity2016pointer for calibration and initial fine-tuning, and SlimPajama cerebras2023slimpajama for extended fine-tuning experiments.
+Calibration: We randomly select 8 examples (2048 sequence length) from WikiText-2 to compute activation statistics for neuron grouping and initial router construction (Eq.
+27).
 We set $K_{a}=10$ for the activation status record, which sounds counter-intuitive but works best in practice.
-For lightweight fine-tuning, we run $1$ epoch using the Adam optimizer [36] with $\beta_{1}$ = 0.9 and $\beta_{2}$ = 0.95.
-We also employ LoRA [37] with a rank of 8 and $lora\_alpha=32$.
-We fine-tune all the models including the baselines with $2,048$ samples.
-We set different initial learning rates for the score scale $\mathbf{u}$ and other parameters, i.e.
-$0.001$ and $5.95e^{-5}$, respectively.
-We set the bias update speed $\gamma$ to $0.001$.
-V-A Main Results
-We compare CMoE with the up-to-date baseline LLaMA-MoE [11], in which the neurons are randomly split, and continual pre-training is carried out with additional router networks.
-The experiments cover both training-free and lightweight fine-tuning versions.
-We design these experiments to demonstrate the remarkable post-training performance of CMoE.
-we demonstrate the results on two different datasets, i.e.
-WikiText-2 [35] and C4 [38].
-The baseline models are chosen as LLaMA-2-7B and LLaMa-3-8B.
-Language Modeling Performance.
-We evaluate the perplexity of the constructed Mixture-of-Experts (MoE) models in Table I.
-We use abbreviations to denote the composition of experts.
-For instance, ‘S2A2E16’ represents 2 shared experts, 2 activated routed experts, and a total of 16 experts and the total activation ratio is $(2+2)/16=25\%$.
-Our findings indicate that in the training-free scenario, the ppl of our baseline LLaMA-MoE on both LLaMA-2-7B and LLaMA-3-8B are NaN across different datasets.
-In contrast, the perplexity of the proposed CMoE can be effectively controlled.
-Moreover, after fine-tuning, we observe that the perplexity can be reduced to as low as 12.73 when the activation ratio is 25%, which corresponds to a sparsity of 75%.
-Additionally, the CMoE model outperforms all state-of-the-art (SOTA) structured pruning methods [9, 22, 23].
-Downstream Tasks Performance.
-We also evaluate the performance of the constructed MoE models on various downstream tasks.
-We evaluate on the following benchmarks: 32-shot BoolQ [39], 0-shot PIQA [40], 0-shot SciQ [41], 5-shot Winogrande [42], 25-shot ARC-Challenge [43], and 10-shot HellaSwag [44].
-The results are presented in table II, where we denote fine-tuning or not with ‘FT’.
-In all activation ratio configurations, we find that the proposed CMoE outperforms the LLaMA-MoE in accuracy across a diverse set of downstream tasks, on both training-free and fine-tuning scenarios.
-To illustrate the performance of the models, we select the SciQ dataset for zero-shot testing and the BoolQ dataset for few-shot testing as representative examples.
-When the activation ratio is set at a fixed value of 25% (corresponding to a sparsity rate of 75%), after fine-tuning, the LLaMA-MoE model can only achieve an accuracy equivalent to 21.2% of that of the relevant dense model on the SciQ dataset and 45.31% on the BoolQ dataset.
-Given that the perplexity (ppl) of the LLaMA-MoE model is Not a Number (NaN) in the training-free scenario, we do not report its accuracy in this case.
-In contrast, even without fine-tuning (i.e., in the training-free scenario), the proposed CMoE model demonstrates superior performance, attaining an accuracy equivalent to 56.3% of that of the dense model on the SciQ dataset and 54% on the BoolQ dataset.
-After fine-tuning, the performance of the CMoE model is further enhanced, with the accuracy reaching 76.59% on the SciQ dataset and 74.3% on the BoolQ dataset (relative to dense model).
-V-B Ablation Studies
-We conduct ablation studies with LLaMA-2-7B and data randomly selected from WikiText-2 training datasets.
-Figure 4: Trade-off between Model Performance and Construction Time with Increasing Training Data.
-TABLE II: Comparison of downstream tasks. Randomly select samples from WikiText-2 training dataset as calibration and training data.
-Model Method Type FT BoolQ(32) SciQ PIQA WinoGrande(5) ARC-C(25) HellaSwag(10) LLaMA-2-7B Dense - - 82.04 90.80 78.78 73.95 53.15 78.55 LLama-MoE A2E8 37.83 20.00 49.73 50.12 25.79 26.18 LLama-MoE A4E16 37.82 20.60 49.56 49.40 25.98 27.21 CMoE S1A1E8 46.09 65.30 52.77 48.70 23.80 30.12 CMoE S1A3E16 50.79 62.00 52.29 50.12 24.23 28.76 CMoE S2A2E16 53.85 66.50 53.26 49.33 23.72 29.85 CMoE S1A1E8 55.04 77.50 57.12 54.06 27.56 38.79 CMoE S1A3E16 53.66 75.30 56.37 53.98 27.82 38.89 CMoE S2A2E16 57.19 77.30 56.86 52.57 26.45 38.90 LLaMA-3-8B Dense - - 83.48 94.2 80.79 77.50 59.72 82.16 LLama-MoE A2E8 37.82 20.30 49.02 50.74 25.68 25.76 LLama-MoE A4E16 37.83 20.50 51.52 49.96 25.27 26.00 CMoE S1A1E8 47.43 47.70 52.56 50.67 23.72 28.08 CMoE S1A3E16 41.31 46.60 51.63 48.77 24.15 27.41 CMoE S2A2E16 45.08 53.10 50.87 53.19 24.23 28.44 CMoE S1A1E8 56.88 72.00 57.34 51.93 25.77 36.68 CMoE S1A3E16 60.55 71.40 57.07 53.51 26.54 36.28 CMoE S2A2E16 62.14 72.20 59.08 51.14 27.73 36.68
-Figure 5: Ablation studies:(a) Impact of shared expert ratio on model performance; (b) Activation Rate vs. Model Performance; (c) Effect of Load Balancing.
-Impact of Training Data Size.
-To quantify the relationship between training data volume, model performance, and computational efficiency, we systematically vary the number of training samples from 0 (initial state) to 4,096 while measuring perplexity (PPL) and construction time.
-Fig.
-4 illustrates the results on the setting of “S2A2E16”, revealing critical trends in performance-cost trade-offs.
-Increasing training data from 0 to 4,096 samples reduces perplexity by 80.4% (62.30 $\to$ 12.21) but non-linearly increases construction time from 298s to 4,502s.
-While performance improves sharply initially (64 samples cut PPL by 63%), gains diminish beyond 1,024 samples (13.47 PPL), with marginal improvements (12.21 PPL at 4,096 samples) requiring disproportionately higher runtime (+133% from 2,048 to 4,096 samples).
-The results demonstrate that CMoE only needs a modest of data to achieve low perplexity, while further performance increase is hard and possibly demands more diverse data.
-Shared Experts Ratio.
-We carry out an extensive analysis of the influence of the shared expert ratio on perplexity.
-In our experimental setup, we have a total of 32 experts, among which 8 are activated.
-We systematically adjust the proportion of shared experts within these 8 activated experts.
-As depicted in Table Fig.
-5, the results clearly demonstrate a distinct trend: as the ratio of shared experts rises from 0.125 to 0.875, the perplexity continuously decreases.
-There is a remarkable improvement in the perplexity value, dropping from 14.48 to 11.93.
-This finding indicates that shared experts generally contribute positively to the performance of the model.
-However, it is also evident that the marginal returns tend to diminish when the ratio of shared experts reaches a relatively high level.
-Activation Rate.
-We conduct an external analysis to investigate how the total activation rate (comprising both shared and activated routed experts) impacts model performance across different domains, in comparison with dense baselines.
-We perform evaluations on the CMoE model featuring a total of 16 experts.
-Specifically, we vary the ratio of shared experts and activated routed experts at a fixed proportion of 1:1 on the WikiText-2 and C4 datasets.
-As illustrated in Table Fig.
-5, a monotonic decrease in PPL can be found as the activation rate increases, gradually approaching the performance level of the dense model.
-For both datasets, an activation rate of 75% enables the model to achieve nearly comparable performance to the dense model.
-On the WikiText-2 dataset, the PPL values are 5.79 for the model under consideration and 5.27 for the dense model; on the C4 dataset, the corresponding values are 11.19 and 7.27.
-These results demonstrate that CMoE architectures can attain a performance quality comparable to that of dense models, even when characterized by relatively high sparsity.
+Lightweight Fine-tuning (2k): We fine-tune using LoRA hu2021lora (rank 8, alpha 32) on 2,048 WikiText-2 samples for 1 epoch.
+The optimizer is Adam kingma2014adam ($\beta_{1}=0.9,\beta_{2}=0.95$).
+We use different learning rates for the router scale parameter (0.001) and other LoRA parameters (5.95e-5).
+The load balancing bias update speed $\gamma=0.001$.
+Extended Fine-tuning (SlimPajama): To assess performance recovery potential, we conduct further fine-tuning on subsets of the diverse SlimPajama cerebras2023slimpajama dataset (up to 1.2B tokens), using similar LoRA settings and run 1 epoch.
+4.1 Main Results
+We compare CMoE against several baselines on PPL, downstream task performance and inference three aspects.
+(1) Dense Models: The original Llama-2 7B and Llama-3 8B checkpoints.
+(2) LLaMA-MoE llama-moe : Random neuron split followed by continual pre-training (we evaluate its training-free state and after our lightweight 2k fine-tune for comparison).
+(3) LLaMA-MoE-v2 llama-moe-v2 : Initialize router networks based on hidden features input and do importance-based neuron partition by measuring the loss change of neuron pruned.
+Our main CMoE configurations vary the activation ratio (Total Activated Experts / Total Experts) while keeping the ratio of shared-to-routed experts 1:1.
+For example, $25\%$ activation uses S1A1E8 (1 shared + 1 active routed / 8 total).
+For baselines, we also use 8 total experts in default.
+Table 1: Language modeling perplexity (PPL) on WikiText-2 and C4 validation sets.
+Lower is better.
+CMoE variants vary activation ratio (fixed 1:1 shared/routed).
+‘TF’: Training-Free, ‘FT’: LoRA fine-tuned on 2k WikiText-2 samples.
+Method Act.
+Llama-2 7B Llama-3 8B WikiText-2 C4 WikiText-2 C4 TF FT TF FT TF FT TF FT Dense 100% 5.27 - 7.27 - 6.14 - 9.44 - LLaMA-MoE 25% nan 468.00 nan 2660.68 nan 988.20 nan 7521.83 LLaMA-MoE-v2 25% >10k 17.72 >10k 55.10 >60k 2754.80 >50k 5300.97 CMoE 25% 62.30 12.73 136.12 32.37 143.38 21.01 284.19 65.57 LLaMA-MoE 75% >60k 322.20 >50k 2061.88 >500k 970.29 >400k 7079.15 LLaMA-MoE-v2 75% >10k 8.68 7358.15 19.76 >10k 1411.23 >10k 2843.39 CMoE 75% 7.02 5.69 10.98 10.21 10.16 8.26 17.35 17.06
+Table 2: Comparison of downstream tasks. Higher is better. Randomly select samples from WikiText-2 training dataset as calibration and training data.
+Model Method Act.
+FT BoolQ(32) SciQ PIQA WinoGrande(5) ARC-C(25) HellaSwag(10) LLaMA-2-7B Dense - - 82.04 90.80 78.78 73.95 53.15 78.55 LLama-MoE 25% 37.83 20.00 49.73 50.12 25.79 26.18 LLama-MoE-v2 25% 51.25 67.00 56.64 52.88 25.68 35.10 CMoE 25% 46.09 65.30 52.77 48.70 23.80 30.12 CMoE 25% 55.04 77.50 57.12 54.06 27.56 38.79 LLama-MoE 75% 37.83 20.00 49.35 50.28 26.37 25.77 LLama-MoE-v2 75% 49.24 79.50 63.55 59.35 34.81 54.89 CMoE 75% 71.68 88.60 73.94 65.90 41.98 69.21 CMoE 75% 75.72 89.60 74.34 67.40 43.94 70.36 LLaMA-3-8B Dense - - 83.48 94.2 80.79 77.50 59.72 82.16 LLama-MoE 25% 37.82 20.30 49.02 50.74 25.68 25.76 LLama-MoE-v2 25% 48.53 45.80 50.87 48.38 24.91 27.31 CMoE 25% 47.43 47.70 52.56 50.67 23.72 28.08 CMoE 25% 56.88 72.00 57.34 51.93 25.77 36.68 LLama-MoE 75% 37.83 20.60 51.31 49.80 26.19 26.08 LLama-MoE-v2 75% 53.30 50.00 50.00 51.38 24.49 28.42 CMoE 75% 58.29 89.90 72.36 65.03 47.52 66.94 CMoE 75% 69.20 88.70 72.80 66.30 48.80 70.63
+Evaluation on PPL.
+Table 1 presents the perplexity values for the WikiText-2 merity2016pointer and C4 raffel2020exploring validation sets.
+In the training-free scenario, CMoE achieves approximately 60 on WikiText-2 when 25% of its components are activated, thereby showcasing its immediate practical utility, but LLaMA-MoE and LLaMA-MoE-v2 both suffer from catastrophic degradation, resulting in perplexity values greater than 20,000 or even NaN without any training.
+When subjected to a lightweight 2k fine-tuning process, CMoE experiences a notable reduction in perplexity across all activation ratios as expected.
+Specifically, 75% activation setting allows CMoE to perform on par with the dense model, evidenced by nearly PPL values.
+Further enhancements through extended fine-tuning is explored in the Ablation Studies section.
+Evaluation on Downstream Task Performance.
+Table 2 presents zero-shot and few-shot results on benchmarks including: 32-shot BoolQ clark2019boolq , 0-shot PIQA bisk2020piqa , 0-shot SciQ welbl2017crowdsourcing , 5-shot Winogrande sakaguchi2021winogrande , 25-shot ARC-Challenge clark2018think , and 10-shot HellaSwag zellers2019hellaswag .
+Similar trends emerge: CMoE is functional training-free, substantially outperforming other methods like LLaMA-MoE and LLaMA-MoE-v2 that collapse without training.
+Lightweight 2k fine-tuning yields significant gains; for the $25\%$ activation CMoE, we recover $>76\%$ of the dense model’s accuracy on SciQ and BoolQ.
+Increasing the activation ratio further closes the gap to the dense model.
+Comparisons with Dense-Small models and the impact of extended fine-tuning are detailed in the Ablation Studies (Section 4.2).
+Evaluation on Inference Speedup.
+We measure inference latency on one single GPU.
+Table 3 shows speedups for the MLP layers and the full model end-to-end.
+Even with the overhead of routing and non-MLP layers, CMoE achieves substantial end-to-end speedups, ranging from 1.4×1.4$\times$1.4 × to 1.6×1.6$\times$1.6 × depending on batch size and configuration.
+This validates that CMoE can achieve practical efficiency enhancements, even though there is a slight performance compromise at high levels of sparsity.
+Table 3: Measured inference speedup for various CMoE activation configurations. S: Shared experts; A: Active routed experts; E: Total experts. E.g. S1A1E8 (1 shared + 1 active routed) / 8 total.
+Configuration Act.
+Memory-Bound Compute-Bound FFN Speedup Full Model Speedup FFN Speedup Full Model Speedup S1A1E8 $25\%$ 3.6×3.6$\times$3.6 × 1.5×1.5$\times$1.5 × 3.75×3.75$\times$3.75 × 1.6×1.6$\times$1.6 × S1A3E16 $25\%$ 3.2×3.2$\times$3.2 × 1.4×1.4$\times$1.4 × 3.45×3.45$\times$3.45 × 1.5×1.5$\times$1.5 × S3A3E8 $75\%$ 1.25×1.25$\times$1.25 × 1.05×1.05$\times$1.05 × 1.27×1.27$\times$1.27 × 1.12×1.12$\times$1.12 × S3A9E16 $75\%$ 1.15×1.15$\times$1.15 × 1.02×1.02$\times$1.02 × 1.09×1.09$\times$1.09 × 1.05×1.05$\times$1.05 ×
+4.2 Ablation Studies
+We conduct ablations using Llama-2 7B and WikiText-2 data unless otherwise noted.
+Impact of Fine-tuning Data Size.
+Figure 2 illustrates the trade-off between fine-tuning data size (using WikiText-2 samples) and construction time versus perplexity (PPL) for the CMoE $25\%$ activation configuration.
+Performance improves rapidly initially when increasing data from 0 to 1,024 samples, but gains diminish significantly afterwards, while the time cost increases non-linearly.
+For instance, increasing samples from 2,048 to 4,096 only marginally improves PPL (12.72 to 12.21) but increases runtime by 133%.
+This suggests that lightweight fine-tuning (e.g., 2k samples) is highly effective, and substantial further gains likely require more diverse datasets like SlimPajama (see Section 4.2) rather than simply more of the same fine-tuning data.
+Figure 2: Trade-off between Model Performance (PPL) and Construction Time with Increasing Fine-tuning Data (WikiText-2 samples, $25\%$ activation).
 Load Balancing.
-CMoE can inherently achieve effective load-balancing among routed experts, with the exception of some special blocks.
-For example, the last block of LLaMA-2-7B exhibits extremely unbalanced expert counts.
-However, this issue can be effectively resolved by the load-balancing mechanism we introduced in Section IV-C.
-As presented in Fig.
-5, prior to the implementation of the load-balancing strategy, substantial disparities in the workloads assigned to different experts are clearly observable.
-Specifically, Expert 3 and Expert 8 are burdened with disproportionately high loads, handling 11,163 and 11,225 instances respectively.
-In stark contrast, other experts process a significantly lower number of instances.
-For example, Expert 12 and Expert 13 handle only 100 and 102 instances respectively.
-Upon the application of our proposed load - balancing mechanism, a more equitable and uniform distribution of the computational workload across all experts is successfully achieved.
-After the adjustment, the number of instances processed by each expert falls within the range of 1,443 to 3,584.
-VI Conclusion
+Figure 3 demonstrates the effectiveness of our load balancing mechanism (Section 3.3) on the final layer of Llama-2 7B, which exhibits highly skewed activation without balancing.
+The mechanism successfully distributes load more evenly across experts.
+Figure 3: Effect of Load Balancing on expert utilization in Llama-2 7B final block ($25\%$ activation).
+Comparison with Dense-Small Models and Extended Fine-tuning.
+Table 4 compares CMoE ($25\%$ activation) against Pythia-1.0B and TinyLlama-1.1B, which have roughly similar active parameter counts.
+The table also shows CMoE performance after fine-tuning on up to 1.2B tokens from the diverse SlimPajama dataset.
+Training-free CMoE underperforms the pre-trained dense-small models.
+However, after lightweight 2k fine-tuning, CMoE becomes competitive.
+With extended fine-tuning (1.2B tokens), CMoE surpasses Pythia-1.0B on several benchmarks and approaches TinyLlama-1.1B, despite using vastly less training data than either (1.2B vs 299B/3T tokens).
+This demonstrates that CMoE effectively transfers knowledge from the base model and benefits significantly from further, albeit still relatively modest, fine-tuning.
+Table 4: Comparison with smaller dense models after fine-tuning on SlimPajama. CMoE uses $25\%$ activation (S1A1E8 configuration).
+Model Training Data PIQA BoolQ HellaSwag Winogrande Pythia-1.0B 299B 69.21 57.83 47.16 53.43 TinyLlama-1.1B 3T 73.29 57.83 59.20 59.12 CMoE-25% training-free 52.77 46.09 28.08 50.67 CMoE-25% 2k WikiText-2 57.12 55.04 36.68 51.93 CMoE-25% 0.5B SlimPajama 59.41 57.13 41.84 53.67 CMoE-25% 1.2B SlimPajama 61.64 59.51 44.02 55.88
+Impact of Expert Configuration ($25\%$ Activation).
+We compare different expert configurations while keeping the total activation ratio fixed at $25\%$.
+Table 5 shows that configurations like S1A1E8 and S2A2E16, which balance shared and routed experts, generally perform well.
+S1A1E8 slightly edges out S2A2E16 in PPL, potentially due to simpler routing, while S2A2E16 shows slightly better performance on BoolQ.
+S1A3E16, with more complex routing (3 active routed experts), shows slightly worse performance.
+Table 5: Performance of different CMoE configurations at a fixed $25\%$ activation ratio (Llama-2 7B).
+Configuration WikiText-2 $\downarrow$ C4 $\downarrow$ BoolQ $\uparrow$ SciQ $\uparrow$ TF FT TF FT TF FT TF FT S1A1E8 60.86 12.76 135.61 32.12 46.09 55.04 65.30 77.50 S2A2E16 62.30 12.73 136.12 32.37 53.85 57.19 66.50 77.30 S1A3E16 89.19 13.84 180.65 33.56 50.79 53.66 62.00 75.30
+5 Conclusion
 We present CMoE, a framework that efficiently carves sparse Mixture-of-Experts (MoE) architectures from dense LLMs through parameter reorganization and lightweight adaptation.
@@
 The router is analytically initialized from activation statistics and refined via differentiable scaling and load balancing.
-Experiments show that CMoE can construct effective Mixture-of-Experts (MoE) models, achieving comparable perplexity to dense models and outperforming baseline models.
-For instance, on the SciQ dataset, CMoE can reach 56.3% even without fine-tuning and it can further improve the accuracy to 76.59% after fine-tuning, a huge outperforming compared with 21.2% (the accuracy of the baseline LLaMA-MoE).
+Experiments demonstrate that increasing the activation ratio to 75% achieves remarkable results, delivering lossless precision from the perpective of the perplexity while still maintaining a 5% acceleration.
+After a brief LoRA fine-tuning process (requiring only 1 hour and 2,000 samples) successfully recovers over 76% of the dense model’s downstream accuracy, highlighting the practical efficiency of this approach for resource-constrained deployment scenarios, which outperforms the baselines LLaMA-MoE and LLaMA-MoE-v2.
 Extensive experiments have demonstrated that CMoE offers a practical approach for deploying large language models (LLMs) in resource-constrained environments.
+Appendix A Algorithmic Analysis
+Figure A.2: The proposed framework, CMoE, and the numerical analysis on its effectiveness.
+As demonstrated in LABEL:fig:hs, the distribution of the FFN hidden state $\mathbf{h}$ is sharply peaked at 00 and constrained within a small range.
+This indicates that most $h_{i}$ are concentrated near zero, confirming the sparsity of activations.
+The symmetry of the distribution suggests a near-normal distribution centered around the mean.
+These observations of high differentiation and independence across different neurons suggest that hidden state values serve as a robust basis for judging neuron activation, motivating the Absolute Top-K (ATopK) metric defined in Equation 4.
+To quantify neuron activation patterns, we first collect activation markers from a calibration dataset.
+Given a batched input tensor $\mathbf{X}\in\mathbb{R}^{b\times s\times d}$, with $b$ the batch size and $s$ the sequence length, we obtain the batched hidden state $\mathbf{H}\in\mathbb{R}^{b\times s\times d_{h}}$ as follows:
+$\mathbf{H}=\text{Swish}(\mathbf{X}\mathbf{W}_{gate})\odot(\mathbf{X}\mathbf{W}_{up}).$ (27)
+Note that in practical implementation, we normalize $\mathbf{X}$, $\mathbf{W}_{gate}$ and $\mathbf{W}_{up}$ before the calculation to eliminate the influence of their magnitudes on the output.
+Using the ATopK metric (defined in the main text), we then calculate the activation markers for the hidden state of all the tokens.
+These markers are reshaped into an activation feature matrix $\mathbf{A}\in\mathbb{R}^{q\times d_{h}}$, where $q=b\cdot s$.
+Denote the $i$-th column of $\mathbf{A}$ as the feature vector $\mathbf{c}_{i}\in\mathbb{R}^{q}$, i.e., $\mathbf{A}=[\mathbf{c}_{1}\;\mathbf{c}_{2}\;\cdots\;\mathbf{c}_{d_{h}}]$, which represents the $i$-th neuron’s activation status on the sampled $q$ tokens.
+By calculating the expected value of each $\mathbf{c}_{i}$, we can obtain the vector of activated rates $\boldsymbol{\mu}$ as:
+$\boldsymbol{\mu}=[\mu_{1},\mu_{2},\cdots,\mu_{d_{h}}],\mbox{ with }\mu_{i}=\frac{1}{q}\sum_{j=1}^{q}c_{j,i},$ (28)
+where $c_{j,i}$ indicates the $j$-th element of $\mathbf{c}_{i}$.
+The histogram of these activation rates $\boldsymbol{\mu}$ is shown in LABEL:fig:rates.
+The histogram reveals a highly skewed distribution of activation rates, where the majority of neurons exhibit low activation rates (below 0.1), with a sharp peak near 0.07.
+However, the distribution also features a long tail, indicating the presence of a subset of neurons with significantly higher activation rates extending up to 1.
+These high-activation neurons are likely active across a wide range of input tokens, making them suitable for processing common knowledge rather than task-specific specialization.
+Therefore, we identify neurons for shared experts by grouping these high-activation neurons.
+Given the total number of shared experts as $N_{s}$ and the expert size $m$, we get the selection indices set $S_{N_{s}}$ by selecting $N_{s}\cdot m$ neurons with highest activation rates based on $\boldsymbol{\mu}$:
+$S_{N_{s}}=\{i\ :\ \mu_{i}\in\text{TopK}(\{\mu_{j}\mid 1\leq j\leq d_{h}\},N_{s}\cdot m)\}.$ (29)
+These indices $S_{N_{s}}$ are then used to form the shared experts by assigning the corresponding parameters from the original FFN, as detailed in Section 3.1 (see Equation LABEL:eq:expert_w in the main text).
+The majority of low activation rates also encourage us to construct routed experts, which are not always activated but are specialized for tokens encountered.
+Appendix B Impact of Calibration Data Size
+We vary the number of calibration samples (WikiText-2) used for neuron analysis.
+Table B.1 shows that performance improves modestly with more calibration data, with diminishing returns beyond 32 or 64 samples.
+Using 8 samples provides a good balance of efficiency and effectiveness.
+Table B.1: Impact of calibration data size on PPL (Llama-2 7B, CMoE $25\%$ activation, training-free).
+Calibration Samples WikiText-2 PPL C4 PPL
+8 60.86 135.61
+32 58.71 129.62
+64 55.88 125.02
+Appendix C Scaling to Larger Models
+To verify that CMoE scales to larger models, we applied the same conversion process (using $25\%$ activation, S1A1E8 configuration) to Llama-2 13B.
+Table C.1 shows the training-free PPL results using different calibration datasets.
+We observe similar patterns to the 7B model: performance improves with more calibration data, and using domain-matched calibration data (C4 for C4 PPL) yields better results.
+This confirms the applicability of CMoE to larger base models.
+Table C.1: CMoE ($25\%$ activation, training-free) performance on Llama-2 13B with varying calibration data.
+Calibration Data WikiText-2 PPL C4 PPL
+8 from WikiText-2 68.82 193.66
+32 from WikiText-2 56.25 167.25
+32 from C4 116.77 134.93
+32 WikiText-2 + 32 C4 68.59 140.17
+Appendix D Discussion
+Limitations While CMoE demonstrates promising results in transforming dense LLMs into efficient sparse MoE architectures, certain aspects merit consideration for future exploration.
+Firstly, the performance of neuron activation profiling and subsequent expert grouping is influenced by the choice of the calibration dataset.
+Optimal results are best achieved when the calibration data aligns well with the characteristics of the target domain or downstream tasks.
+Secondly, while the training-free router construction provides notable computational efficiency, it may not fully capture all nuances of routing decisions in the same way a router trained extensively end-to-end with the experts might.
+In some scenarios, particularly those requiring highly dynamic expert selection, there might be opportunities for further performance gains through additional fine-tuning.
+Future work could also explore the application and adaptation of CMoE across a wider array of model architectures and a more diverse set of downstream tasks to further understand its broader applicability and potential refinements.
+Impact Statement This paper aims to contribute to the advancement of the Machine Learning field.
+While our work may have various societal implications, we do not find it necessary to emphasize any particular consequences here.